### v1

# From #Dr00gtiktok to #harmreduction: Exploring Substance Use Hashtags on TikTok

Abstract
The rise of TikTok as a primary source of information for youth, combined with its unique short-form video format, creates urgent questions about how substance use content manifests and spreads on the platform. This paper provides the first in-depth exploration of substance use-related content on TikTok, covering all major substance categories as classified by the Drug Enforcement Agency. Through social network analysis and qualitative coding, we examined more than 2,333 hashtags across 39,509 videos, identified 16 distinct hashtag communities and analyzed their interconnections and thematic content. Our analysis revealed a highly interconnected small-world network where recovery-focused hashtags like “#addiction”, “#recovery”, and “#sober” serve as central bridges between communities. Through manual coding of 351 representative videos, we found that Recovery Advocacy content (33.9%) and Satirical content (28.2%) dominate, while direct substance depiction appears in only 26% of videos, with active use shown in just 6.5% of them. This suggests TikTok functions primarily as a recovery support platform rather than a space promoting substance use. We found strong alignment between hashtag communities and video content, indicating organic community formation rather than attempts to evade content moderation. Our findings inform how platforms can balance content moderation with preserving valuable recovery support communities, while also providing insights for the design of social media-based recovery interventions.
Introduction
Social media platforms have become the primary lens through which young people view and understand the world, with TikTok emerging as a dominant force among youth - reaching 63% of teens, with 17% reporting near-constant use (Vogels, Gelles-Watnick, and Anderson 2023). More concerning is that TikTok has become a primary news source for this demographic, often replacing Google for information-seeking (Walker and Matsa 2023).
Figure 1: Examples of TikTok videos that can be found in the substance use communities based on hashtag use
This shift in information consumption patterns takes on particular urgency when considering content related to illicit substances. While overall drug use among 8th, 10th, and 12th graders showed promising declines during COVID-19 social distancing (National Institute on Drug Abuse 2022), this drop in usage may have masked how drug-related discussions were shifting online. Social media platforms can rapidly disseminate trends and behaviors through psychological processes like social comparison and fear of missing out (FOMO) (Montag, Yang, and Elhai 2021), with adolescents often imitating influencers and peers they follow. Although most social media platforms use content moderation, their rapid and viral nature can make it challenging to effectively regulate exposure to harmful content (Boyd 2014; Lookingbill and Le 2024), potentially amplifying dangerous trends. Therefore, understanding such emerging trends on social media platforms such as TikTok, Instagram, and Reddit (Rutherford et al. 2022a, 2023; Tan and Weinreich 2021; Bouzoubaa, Young, and Rezapour 2024) becomes more critical. For instance, the “xandemic” between 2016-2018, characterized by high school students documenting the recreational use of Xanax to the point of blackout, exemplifies how social media trends can blur the lines between casual and harmful behaviors and even normalize dangerous substance use behaviors (Fuller et al. 2024). Similarly, platform-specific hashtag “#pingtok” - a portmanteau of the slang term for an MDMA pill, pinger, and TikTok - which has accumulated over 26,000 posts at the time of writing, demonstrates how substance-related content can become embedded in platform culture, with its own specific terminology and community. These trends are especially concerning given robust evidence that early exposure to substance use significantly increases the risk of developing substance use disorders (SUD) in adulthood (McCabe et al. 2022; Volkow and Wargo 2022; Odgers et al. 2008).
Previous research on substance-related content on TikTok has typically focused on either specific substances such as cannabis or nicotine (Rutherford et al. 2023, 2022a; Morales, Fahrion, and Watkins 2022; Tan and Weinreich 2021) or isolated trending phenomenon like “#pingtok” (Whelan, Noller, and Ward 2024). While some researchers have examined hashtag networks and content moderation strategies around other sensitive topics like eating disorders (Bickham et al. 2024) and how hashtags create communities around shared experiences (Eriksson Krutrök 2021), there remains limited understanding of how hashtags shape substance-related discourse on TikTok. Moreover, no studies have comprehensively examined how substance-related content manifests across TikTok’s ecosystem of communities, particularly through its short-form video format and hashtag-driven discovery system (Abidin 2020). To address this gap, we draw from the Drug Enforcement Agency’s comprehensive drug classification system and examine content from every major substance category using a mixed-methods investigation of TikTok content. Our research questions are:
  • RQ1: What is the network structure of substance-related hashtags on TikTok, and what distinct communities emerge within this network?
  • RQ2: What are the topics and key characteristics of the videos associated with the hashtags in these communities?
  • RQ3: What are the effects of substance visibility on topic distribution and engagement metrics?
For this purpose, we developed a list of 585 seed hashtags related to substance use and collected more than a thousand videos for each of them, resulting in a dataset of 504,294 hashtags and 272,910 videos  [^1] Datasets, including seed hashtags, and code are available at https://github.com/social-nlp-lab/explore˙tiktok˙su. Videos are not directly provided but can be accessed via URLs.. After filtering relevant hashtags, we identified 2,333 exemplary hashtags covering the broad categories of substance use, and 351 videos for qualitative coding. Using a combination of network analysis of hashtag co-occurrence and in-depth content analysis, our analysis showed a highly interconnected small-world network where recovery-focused hashtags like “#addiction”, “#recovery”, and “#sober” serve as central bridges between communities. Within these communities, we found that Recovery Advocacy content (33.9%) and Satirical content (28.2%) dominate, while direct substance depiction appeared in only 26% of videos, with active use shown in just 6.5%. This suggests TikTok functions primarily as a recovery support platform rather than a space promoting substance use.
Through this work, we contribute to social computing and substance use literature by providing the first comprehensive mapping of substance-related content on TikTok. Our findings can inform platform policies around content moderation strategies that preserve supportive recovery communities while better identifying potentially harmful content, guide the development of social media-based recovery support interventions, and help design youth substance use prevention efforts that account for how substance-related content actually manifests on TikTok.
Related Works
As digital platforms become ubiquitous, understanding their impact, especially on young people’s health behaviors regarding substance use, is more crucial than ever. On the one hand, social media emerged as a medium for facilitating sensitive discussions and disclosure of lived experiences, including mental health and eating disorders (Rutherford et al. 2022a; Bickham et al. 2024; Bouzoubaa et al. 2024). On the other hand, several studies have examined the relationship between online peer groups, social media homophily, and substance use, finding positive correlations between substance use and social media engagement. For example, Willoughby et al. (2024) found that young adults reported an increased intention to use cannabis after exposure to pro-cannabis messages online. Similarly, Sun et al. (2023) showed that exposure to e-cigarette advertisements on social media platforms is associated with adolescents’ e-cigarette usage in later years. Miller et al. (2021) examined how online and offline peer interactions influence substance use behaviors and found that online belonging and social media homophily predicted regular stimulant and opioid use among substance users. Hanson et al. (2013) have studied online discussions related to stimulant drug use on Twitter and found a high correlation between the popularity of such topics and final exam periods. Geusens et al. (2023) found that higher levels of alcohol and marijuana consumption predicted later social media posts about those substances, but posting did not predict later substance use. Sun et al. (2023) found a positive correlation between early exposure to e-cigarette advertisement and subsequent lifetime usage.
Hashtags, a medium for interactions on platforms like Facebook or Twitter, have been studied comprehensively in the past (Mousavi and Ouyang 2021; Ince, Rojas, and Davis 2017; Small 2011; Rezapour et al. 2017). People can contribute to the narrative by hash-tagging their own personal experiences and thoughts (Yang 2016). At their core, hashtags facilitate content discovery and community creation by allowing users to contribute to the discussion through keyword tagging (Small 2011). Previous studies have examined substance use discourse across various social media platforms, contributing to our understanding of how substance use behaviors manifest online. For instance, on Instagram, investigating “#studydrugs” demonstrated a predominantly positive portrayal of stimulant use for academic performance enhancement (Petersen et al. 2021).
On TikTok, content creators deliberately employ hashtags like “#fyp” (For You Page) to increase visibility and maximize audience reach within TikTok’s recommendation system (Abidin 2020). The platform’s ranking algorithms connect users based on hashtag engagement patterns, creating what Eriksson Krutrök (2021) calls “algorithmic closeness” - subcommunities united through shared content interests. This algorithmic dynamic shapes how substance-related discourse spreads and evolves on TikTok. For example, research examining TikTok hashtags like “#puffbar” has revealed concerning patterns of youth-oriented substance marketing and normalization (Tan and Weinreich 2021), while analysis of “#stonertok” and “#420vibes” showed the majority of cannabis-related videos positively portray substances (Rutherford et al. 2022b).
Our study advances existing research by employing a network-based and qualitative approach to explore the interactions and communities associated with substance-related hashtags on TikTok. While previous research has thoroughly documented individual communities, the dynamics of how these groups interact across the platform’s broader ecosystem remain understudied. Our work can help social media platforms like TikTok refine their policies to better distinguish between harmful content requiring removal and harm reduction information that should be preserved, contributing to more effective content moderation.
Methods
Data
Seed Hashtag Selection: To develop an initial set of seed hashtags related to substance use for the data collection, we first identified hashtags previously used in studies of substance-related content on social media platforms (Petersen et al. 2021; Russell et al. 2022; Rutherford et al. 2023, 2022b; Singh and Wu 2021) as well as searching for hashtags on the platform associated with substance names (both formal and slang) listed on the Drug Enforcement Agency’s Fact Sheet (United States Drug Enforcement Administration 2025). Using this set, we conducted iterative searches on TikTok to identify additional relevant hashtags by examining co-occurring hashtags within the content. This process continued until we reached saturation, resulting in a final set of 585 substance-use-related hashtags. [^2] The list will be shared upon the paper’s acceptance.
Data Collection: Using the Unofficial TikTok API [^3] https://github.com/davidteather/TikTok-Api, we collected data for around 1,000 algorithmically served videos for each of the 585 seed hashtags. This resulted in a dataset of 275,291 videos and a total of 504,294 unique hashtags.
Figure 2: Data selection and pre-processing pipeline to create the filtered set.
Data Preprocessing and Network Refinement: To enhance the feasibility of our analysis and reduce noise from potentially irrelevant content, we filtered the data to include only videos that featured at least two of the seed hashtags. This filtering step resulted in a refined dataset of 39,602 videos and 99,727 unique hashtags (Figure 2). To further clean the data, we constructed a co-occurrence network where nodes represent individual hashtags and edges represent their co-occurrence in videos. More specifically, edge weights were defined by the frequency with which two hashtags appeared together in videos. We then employed a two-stage filtering process to focus on the most relevant and interconnected hashtags. First, we removed hashtags with low connectivity by filtering out nodes in the bottom 25th percentile of degree centrality. This step eliminated peripheral hashtags with minimal connections to the broader substance use discourse, removing those that were incidentally rather than substantively related to substance use discussions. Next, we applied k-core decomposition ($k=10$) to identify the densely connected core of the network (Seidman and Foster 1978). K-core decomposition recursively removes nodes with degrees less than $k$ until all remaining nodes have at least $k$ connections, ensuring a cohesive network structure. This threshold was chosen based on exploratory analysis, which revealed that $k=10$ produced the most structurally cohesive subset of hashtags while retaining a sufficient number of nodes and edges to analyze the discourse. The resulting refined network of hashtags contained 2,333 nodes and 46,990 edges, representing the most interconnected hashtags in substance use-related content.
Communities Definition Example Hashtags
Emotions and Feelings Words and phrases related to emotional states and feelings. #love, #gratitude
Health Conditions Words and phrases related to health issues directly related to substance use, addiction-related health problems, or chronic conditions that may lead to substance use. #addiction, #chronicpain
Alcohol Words related to alcoholic beverages and spirits. This is just about the substances or drinks and not the side effects and consumption methods. #alcohol, #vodka
Cannabis Substances related to marijuana, including recreational and medicinal use. #cannabis, #stonersoftiktok
Cognitive Enhancement Substances about nootropics, smart drugs, and methods to improve cognitive function. #modafinil, #smartdrugs
Commonly-Misused Substances Substances frequently used and misused, such as licit and illicit substances. #heroin, #xans
Consumption Method Words and phrases that show specific ways in which substances (including alcohol, illicit or licit drugs, tobacco, and cannabis) are ingested, administered, or used. #smoking, #injection
Awareness and Advocacy Words related to information and strategies to prevent and raise awareness of substance abuse or reduce associated harm, discussions about drug laws, policies, and related social issues, and content about addiction recovery, sobriety, and support systems. #harmreduction, #recoverylife
Other Substances Mentions of less common substances, oftentimes legal, over-the-counter medications, herbal remedies, or supplements. #vitamins, #magnesium
Platform Tags and features specific to social media engagement, visibility, and trending tactics. #fyp, #viral
Substance Effects Words and phrases that describe the physical or mental effects of substance or alcohol use, both desired effects and side effects. #stoned, #drunk
Tobacco Nicotine Words related to tobacco products, cigarettes, vaping, and nicotine use. #vaping, #ecigs
Humor Words, phrases, or hashtags related to jokes, memes, or any content meant to be funny but specific to substance use, addiction, or recovery. #soberhumor, #drughumor
Location Words related to geographical locations, including cities, states, countries, or continents. #kensingtonphilly, #boston
Occupation Words related to occupations or professions. #nurselife, #medstudent
Identity and Community Hashtags related to any social identity, demographic group, or community affiliation, such as race, ethnicity, gender identity, sexual orientation, disability status, socioeconomic background, immigration status, religion, age group, or membership in specific subcultures or communities. #lgbtqia, #transgender
Misc Any tag that does not fit into the above categories. #cookies, #foundersday
Table 1: Categories for semantically classifying hashtags from TikTok content, their definitions, and example hashtags.
Substance Use Hashtags Network
Community Detection: We first applied established community detection algorithms including Clauset-Newman-Moore (Clauset, Newman, and Moore 2004) and Louvain method (Blondel et al. 2008), but these approaches did not yield clusters that meaningfully captured substance-related discussions. Therefore, we used a hybrid approach combining manual and automated methods to identify thematic groups within the hashtags. We first manually annotated a sample of 251 hashtags and iteratively established thematic categories. As a result, we identified 16 categories representing topics related to substance use and consumption, identity, and community, health and emotion as well as awareness and advocacy. In addition, we created a Misc category for general hashtags, not related to substance use content. The final set of categories and their definitions are shown in Table 1. These categories were then used to prompt GPT-4o (OpenAI 2024) to classify hashtags. When evaluated against human-labeled data, our best-performing prompt achieved an accuracy of $89.36\%$, demonstrating strong reliability in the thematic categorization. We then used this prompt to classify the full set of 2,333 hashtags. The resulting labeled hashtags were mapped back onto the network, creating themed communities that reflect different aspects of substance use discourse on TikTok.
Network Analysis: For both the overall network and each community, we calculated standard network centrality measures including betweenness (Freeman 1977), closeness (Sabidussi 1966), degree (Freeman 1978), and eigenvector centralities (Bonacich 1972). These measures help identify influential hashtags and understand the structure of discourse within and across themes. Additionally, we computed network-level metrics such as average degree, clustering coefficient, and average shortest path length to characterize the overall connectivity patterns and community structure of substance use-related discourse on TikTok.
Characterizing Video Content
Codebook Development: To establish the codebook for coding substance-related TikTok videos, two researchers independently open-coded 160 randomly selected videos. Through this initial coding process, researchers identified nine distinct video topics and coded for substance presence and usage. The physical presence of drugs and related substances was of interest to us in determining which communities and topics were more engaged in depicting substances. The intercoder reliability was strong for coding substance visibility ($\kappa=0.87$) but only fair for video topics ($\kappa=0.35$). After discussing initial disagreements, the topics were refined from nine to seven distinct categories. The coders then had an iterative process for discussing the codes with all researchers on the team to identify coherent topics of videos within the TikTok dataset, resolve any disagreements, and finalize the codebook. The final codebook captured two key dimensions: video topic and substance presence, as shown in Table 3.
Video Selection: We started by selecting all videos that contained at least one of the 2,333 core hashtags, resulting in a dataset of 39,509 videos. To refine this dataset for detailed coding, we identified the top five hashtags per community based on their degree centrality.For each of the 80 selected hashtags, we then filtered videos using a stepwise approach:
  1. 1.
    50 initial videos were acquired from our dataset for each of the 80 hashtags. Ten of each were acquired from each quintile of viewership for the hashtag, achieving representation regardless of virality.
  2. 2.
    Each video was evaluated against three key criteria:
    • Availability: The video is accessible and not moderated or removed by TikTok.
    • Language: The video features English as the primary spoken or written language.
    • Substance Relevance: The video clearly depicts substance-related content, whether through visual cues, discussion, or behaviors.
We continued this process until we identified five videos per hashtag, one per each quintile of engagement, that met all three criteria. In total 977 links were coded, with 852 videos available (87.2%), of those 695 videos were in English (81.5%), and of those 351 videos were substance-related (50.5%). We conducted an inductive content analysis (Elo and Kyngäs 2008) on these 351 selected videos based on the codebook to ensure our final dataset was representative of the substance-use discourse across different communities (See Table 2).
Findings
Filtered Coded
N Videos 39,509 351
Avg. Likes 66.5K 30.9K
Avg. Comments 702.6 305.9
Avg. Shares 3513.2 2652.5
Min. Date 2015-09-22 2019-11-13
Max. Date 2024-05-13 2024-05-09
Table 2: Video statistics of the dataset filtered for RQ1 and the sub-set used for coding in RQ2.
RQ1: Network Structure and Communities
Analysis of hashtags network showed complex, interconnected structures with distinct communities. The overall network contained 2,333 nodes (hashtags) and 46,990 edges (Table C in Appendix), with an average degree of 40.28 and a weighted degree of 513.84 as listed in Table C. These metrics indicate substantial interconnection - each hashtag co-occurs with approximately 40 other hashtags, and when hashtags do co-occur, they appear together frequently (over 500 times on average). The network exhibits small-world properties characterized by a high average clustering coefficient (0.52) coupled with a short average path length (2.20) (Watts and Strogatz 1998). This structure suggests that while substance-related discussions form tight communities, information can still flow efficiently across the network through just two intermediate hashtags on average.
Network centrality analyses revealed two distinct categories of influential nodes. The first category consists of substance-specific hashtags like “#addiction”, “#recovery”, “#sober”, “#wedorecover”, and “#harmreduction”, which scored highly across all centrality measures (betweenness, closeness, degree, and eigenvector). These hashtags function as critical bridges between different substance-related discussions, facilitating information flow between harm reduction, recovery, and other communities. The second category includes general-purpose hashtags like “#fyp” and “#mentalhealth”, which exhibit high centrality but serve primarily to connect substance-related discourse to broader TikTok conversations. This integration with platform-wide hashtags means substance-related content can potentially reach users not explicitly seeking such content through TikTok’s recommendation system.
Our community detection analysis identified 16 different themes (excluding Misc, see Table 1), each revealing distinct patterns in how substance-related content is organized on TikTok (Figure 6 in Appendix). The largest community in this network represents Awareness and Advocacy, containing approximately 12% of all hashtags (280 nodes) and featuring highly central nodes focused on harm reduction (e.g., “#harmreduction”, “#harmreductionworks”), recovery and sobriety (e.g., “#wedorecover”, “#sober”), and public health awareness (e.g., #opioidawareness, #opioidcrisis). The Emotions and Feelings cluster, while smaller with 180 nodes, shows high internal connectivity (clustering coefficient of 0.61) and frequently bridges to other communities through hashtags like “#mentalhealth” and “#anxiety,” suggesting the important role of emotional support in substance-related discussions. The network also revealed unique patterns in how substance-specific discussions are organized. Communities such as Cannabis (142 nodes) and Tobacco/Nicotine (138 nodes), exhibit strong internal clustering but limited external connections. This isolation suggests the discussions stay within their topical boundaries. The Commonly-Misused Substances cluster, though relatively small (120 nodes), contains hashtags often straddling the line between awareness and potential promotion of substance use (e.g., ”#studydrugs”, “#pingertok”, “#smartdrugs”). The Platform-specific cluster (165 nodes) showed distinct characteristics (e.g., “#fyp”, “#viral”), with the highest average betweenness centrality, indicating its crucial role in amplifying substance-related content to broader audiences.
Figure 3: Engagement metrics across select popular hashtags in the coded set. Note, #harmreductionsaveslives has been shorted to HSL.
RQ2: Topics and Characteristics of Videos
Analysis of 351 videos based on our codebook (Table 3) showed that the most prevalent type of content in the videos focused on advocacy and recovery support (33.9%, 119 videos), where creators shared personal recovery journeys, celebrated sobriety milestones, or provided support to others struggling with substance use. For example, one video shows a creator celebrating “100 days sober milestone and encouraging others to seek help.” Satirical or humorous content formed the second largest category (28.2%, 99 videos), with videos often using comedy to address serious topics around substance use. The content is typically presented in the form of acted-out skits or parodies designed to gain engagement. For instance, one video presented a “comedic skit about stereotypical behavior while intoxicated.”
Informational content, typically from claimed medical or professional perspectives, comprised 13.1% (n=46) videos. These videos often featured “doctors explaining the effects of new medications on brain chemistry” or provided educational content about substance effects and risks. A smaller portion focused purely on social media engagement (3.9%, n=14 videos), with creators recording substance-related content primarily to gain followers or views. A sizable portion of the content is in the promotional category, appearing in 7% (n=24) videos. The majority of them portray a demonstration of a vaping device, Juul or lighter, and a few of them promote nootropics or study drugs. Notably, while discussions of substances were common, direct depiction of substances was less frequent - appearing in 25.9% (n=91) of videos, with only 23 of those videos showing active substance use. This suggests a tendency toward discussing rather than displaying substance use.
Dimension Category Definition N (%) Example
Visible Substance Shown Direct visual depiction of drugs or other substances 91(25.9%),with 23(25%) Actively Using Displaying empty prescription medication bottles on a table beside the subject while they reminisce about prior SUD
Not Shown No direct visual depiction of substances 260 (74.1%) discussing recovery journey without showing any substances
Recovery Advocacy Content promoting or raising awareness about recovery efforts 119 (33.9%) Person sharing their 100 days sober milestone and encouraging others to seek help
Satirical Humorous or satirical content about substance use without meaningful substance-related message 99 (28.2%) Comedic skit about stereotypical behavior while intoxicated
Video Topic Informational Educational content about substances from claimed professional/medical perspective 46 (13.1%) Doctor explaining the effects of a new medication on brain chemistry
Promotional Content meant to promote or sell substance-related products 24 (6.8%) Video advertising nicotine vaping products with links to purchase
Social Documentation Content created primarily for social media engagement, focusing on filming substance use 14 (4.0%) Creator recording themselves taking shots at a party while encouraging viewers to like/follow
Trip Reports First-person accounts of substance use experiences 4 (1.1%) Individual describing their experience with a specific substance and its effects
Other Substance-related content that doesn’t fit other categories 45 (12.8%) News clip about substance policy changes
Table 3: Codebook for qualitative analysis of substance-related TikTok videos (N=351)
Figure 4: Sub-networks of Awareness & Advocacy (Pink) and Health Conditions (Blue) to visualize strong overlap between hashtag usage.
Alignment Between Hashtags and Video Content We also examined whether videos’ actual content aligns with their associated hashtag and communities. While each video was coded for its primary topic, we find that videos typically contain hashtags from multiple communities, revealing complex patterns in how substance-related content is tagged and disseminated.
Advocacy, Recovery, and Support Content Videos coded primarily as advocacy and recovery content (119 videos, 33.9%), the largest thematic group in our manual coding, showed strong alignment with two major hashtag communities (see Table 4):
  • Awareness & Advocacy (99 videos, 57.2%): Recovery-focused videos frequently used hashtags like “#harmreduction” and “#wedorecover”. Notably, while recovery-focused hashtags were more commonly used overall, content tagged with harm reduction messaging (particularly “#harmreductionsaveslives”) generated substantially higher user engagement across all metrics - views, likes, comments, and shares (Figure 3).
  • Health Conditions (100 videos, 56.8%): Videos discussing recovery often included medical or health-focused tags like “#addiction”, “#chronicpain”, and “#pain” (Figure 4)
Content of videos seen in this group are typically centered around active or past efforts to recover from substance abuse, discussed in context of cautionary tales, or implicitly imploring the viewer to regard substance abuse with regard and care due to potential harms.
Entertainment and Platform Engagement Entertainment and humor-focused content (99 videos, 28.2%) as well as videos that documented using a substances for engagement (14 videos, 4%) similarly bridged multiple communities. These videos predominantly contained hashtags from the communities:
  • Platform (82 videos, 40.6% of all coded videos): Entertainment videos heavily utilized platform-specific tags like “#fyp” and “#pingtok”
  • Humor (60 videos, 78.9%): Videos using “#addictionhumor” and similar tags often presented substance use through comedy or satire
  • Substance Effects (35 videos, 53.8%): Videos that were satirical or relatable often documented an effect of substance use (e.g., what it’s like being “#high” or “#stoned”). The subjects in these videos may not necessarily be intoxicated but could be enacting or portraying the effects.
Content in this group still discussed topics of substance use and misuse, however the seeming salient topic was humor. A large proportion of videos in this group were meme-based videos, where the original content was either not drug related but brought into drug discussion with filters, captions, or additional context, or was drug related but leaned towards humor and not object lessons.
Figure 5: Distribution of video topics across three levels of substance visibility (no substances shown, substances shown but not used, and active substance use) in TikTok videos.
Educational and Informational Content Videos coded as informational (46 videos, 13.1%) showed connections to multiple professional and health-focused communities:
  • Health Conditions (20 videos, 11.4%): Medical information and health impacts
  • Awareness & Advocacy (29 videos, 16.8%): Harm reduction education
  • Other Substances (16 videos, 24.6%) :
  • Occupation (10 videos, 19.2%): Content from healthcare workers and other professions
Content in this group was thematically clinical or information focused. Humor was used sparingly if at all. The vast majority of videos in this cluster used detailed captions to go over specific facts regarding drug use or were more long-form and had a health professional or recovery advocate speak at length on the topic. Recovery organizations were seen multiple times in the group, with high production value videos.
While advocacy, entertainment, and educational content showed strong alignment with specific hashtag communities, certain clusters served as contextual bridges across all content types. Hashtags from the Commonly Misused Substances (111 videos, 31.6%), Identity & Community (n = 95 videos, 27.1 %), and Occupation (53 videos, 15.1 %) clusters appeared across all seven content categories, suggesting their role in providing context rather than signaling specific content types. For instance, videos tagged with “#opiates” spanned multiple content categories: healthcare workers explaining naloxone administration (Recovery/Advocacy), recovery nurses humorously discussing their reliance on caffeine (Satirical), and medical professionals explaining medication effects (Informational). Similarly, occupation-related hashtags like “#nurse” or “#doctor” and identity-based tags like “#disabled” or “#veteran” appeared across all content. Consumption Method hashtags like “#vape” or “#smoke” (101 videos, 28.8% of videos) served a more specific bridging function, describing substance use behaviors across content types.
RQ3: Effects of Substance Visibility
Given these patterns in how content creators use hashtags to frame and contextualize their videos, we next examined a key aspect of substance-related content: the inclusion of substances or their related paraphernalia themselves in the videos with respect to context and engagement. This analysis revealed distinct patterns in content distribution across three categories: no substances shown (n=260), substances shown but not used (n=68), and active substance use (n=23). Videos showing active substance use were predominantly focused on documentation for social media (approximately 40% of videos, Figure 5), with a stronger tendency towards lifestyle or vlog-style content. In contrast, recovery advocacy content was most prevalent in videos without substance visibility, while promotional content appeared most frequently in videos that showed but did not depict active use of substances (e.g., videos promoting nicotine products). Within videos showing substances, content predominantly featured alcohol, nicotine, or cannabis products, with only one video collage containing “hard drugs” such as Percocet, Adderall, and Xanax. A Kruskal-Wallis H (Kruskal and Wallis 1952) test was performed to compare the means in engagement metrics (views, likes, comments, and shares) between the substance presence categories. Notably, there were no statistically significant differences between groups. This suggests that user engagement on TikTok is driven more by content type and presentation than by the mere presence or use of substances.
Discussion
Social media’s impact on substance use behaviors has been a growing concern, with platforms often criticized for normalizing or promoting dangerous behaviors (Curtis et al. 2018; Moreno and Whitehill 2014; Lobstein et al. 2017; Rutherford et al. 2023; Romer and Moreno 2017). Our analysis of TikTok, however, revealed a more nuanced reality where recovery and harm reduction messaging dominate substance-related discourse. We found evidence of TikTok emerging as a significant platform for recovery support while maintaining its core identity as an entertainment platform.
The Power of Digital Recovery Narratives The most striking finding is the prevalence of recovery-focused content, which typically feature first-person narratives celebrating sobriety milestones or sharing recovery journeys. As Slade et al. (2021) note, such “digital recovery narratives” can increase knowledge of evidence-based treatment options, normalize addiction experiences, and encourage help-seeking behaviors. The impact of these narratives appears to be bidirectional - benefiting creators and viewers. For creators, sharing recovery stories can reinforce recovery-focused, collective identities and provide accountability through community support (Beck et al. 2023). For viewers, these narratives offer realistic models of recovery success, reduce feelings of isolation, and demonstrate that recovery is possible across different circumstances (Jacobson 2001; Kaiser et al. 2020; Corvini, Golomski, and Burns 2024).
Balancing Entertainment and Support in Health Communication The significant presence of humorous content (28.2%) alongside serious recovery narratives reflects TikTok’s unique ability to facilitate health communication through multiple approaches. Rather than undermining serious messages, humor appears to serve as an additional entry point for engaging with recovery-related content. The use of humor in substance-related content may serve important psychological functions, as research has shown that comedy can help individuals process difficult experiences (Martin 2002, 2003) and reduce stigma around mental health challenges (Corrigan et al. 2014). In the context of substance use and recovery on TikTok, videos that integrate humor appear to make the challenges around substance use more approachable while facilitating community connection through shared experiences. This balance between entertainment and support may be particularly effective for reaching younger audiences. However, this dual nature also presents challenges. While most humorous content avoids glorifying substance use some videos may inadvertently normalize concerning behaviors through comedy. This tension between engagement and responsible messaging represents a key challenge for platforms hosting health-related content.
Network Structure Facilitates Community Building The network analysis revealed a highly interconnected structure with small-world properties, enabling both focused discussions within communities and efficient information flow between them. Hashtags serve multiple functions: some act as content summarizers (e.g., “#harmreduction”, “#recovery”), while others serve as contextualizers (e.g., identity-based tags, occupation tags). Within this network structure, we observed the presence of algospeak in the form of coded or evasive hashtags (e.g., “#drvg”, “#drvgzx”, “#dr00gtiktok”) that are designed to evade content moderation. Our content analysis revealed that these coded hashtags were predominantly associated with satirical and relatable content, with a small subset appearing in recovery advocacy videos. This pattern suggests these alternative spellings may serve less to promote substance use and more to facilitate humor-based discussions and shared experiences around substance use, though this practice could complicate platforms’ efforts to effectively moderate substance-related content (Klug, Steen, and Yurechko 2023; Levine 2022). This dual functionality helps explain why content tagged with harm reduction and recovery-focused hashtags received substantially higher engagement despite appearing less frequently in our sample.
Looking Forward: Implications for Social Media Health Communication
Our findings have important implications for how social media platforms approach substance-related content, not just on TikTok but across all platforms featuring short-form video content like Instagram Reels and YouTube Shorts. The success of recovery narratives suggests that brief, personal videos may be particularly effective for health communication. Rather than focusing solely on restricting content, platforms might consider ways to actively promote recovery and harm reduction content, given its demonstrated appeal to users.
The emergence of TikTok as a recovery support space (Russell et al. 2022) offers valuable lessons for content moderation and platform design. While current approaches appear to effectively limit harmful content while allowing space for support and advocacy, refinements could address the challenges of moderating humor-based content. As short-form video becomes an increasingly dominant mode of social media communication, particularly among younger users, understanding how to foster supportive recovery communities while mitigating potential risks becomes important. Several areas warrant further research. While we found evidence of TikTok serving as a positive space for recovery support, there is a pressing need to understand the real-world impacts of this online discourse. Future studies should examine how exposure to recovery content affects help-seeking behaviors and treatment outcomes. Additionally, researchers should investigate potential risks and benefits for people who use drugs who engage with or create TikTok content. This is particularly important for content creators who publicly identify as being in active recovery, as sharing personal narratives may have complex psychological and social implications that are not yet well understood (Rennick-Egglestone et al. ). Additionally, investigating the intersection of online and offline support networks could provide insights into how digital recovery communities complement traditional recovery support services, optimizing online and in-person interventions. This work would be particularly valuable for reaching younger populations, who may be more likely to seek initial support through social media platforms.
Conclusion
This study explored the use of over 2.3K substance-related hashtags used on the popular short-form social media platform, TikTok. Through a semi-automatic community detection process, we identified 16 distinct hashtag communities, including Advocacy and Awareness, Health Conditions, and Humor. Our social network analysis revealed a highly interconnected small-world network (clustering coefficient 0.52, average path length 2.20) with influential hashtags like “#addiction”, “#recovery”, and “#sober” serving as bridges between communities. Through manual coding of 351 representative videos, we found that Recovery Advocacy content dominated (33.9% of videos), followed by Satirical content (28.2%), suggesting TikTok serves primarily as a platform for support rather than substance promotion. This is further supported by our finding that substances were present in only 91 videos (26%), with active use shown in just 23 videos. The alignment between hashtag communities and video content, particularly in recovery-focused videos, indicates organic community formation rather than attempts to evade content moderation. These findings have important implications for platform policies and interventions - while the prevalence of recovery-focused content suggests TikTok may serve as a valuable space for support and advocacy, the significant presence of humorous content alongside serious advocacy highlights the need for nuanced content moderation approaches. Future research should examine how these digital recovery communities complement traditional recovery support services and their real-world impacts on help-seeking behaviors and treatment outcomes, particularly among youth.
Limitations
Several limitations of our study warrant discussion. First, our findings may not fully represent the entire TikTok ecosystem, as they are influenced by limitations in data availability and collection processes. TikTok’s (unofficial) API serves content through its recommendation algorithm, which means we received videos that did not necessarily contain our queried hashtags. While we attempted to address this by sampling videos across engagement quartiles to ensure representativeness, the algorithm’s opacity makes it difficult to guarantee complete coverage of substance-related content. This algorithmic bias may have also led to over-representation of certain creators whose content is favored by TikTok’s recommendation system. TikTok, in our experience, also cuts off data requests arbitrarily. Our approach of using 1,000 videos per hashtag was due to an alignment with the average number of videos we could reliably retrieve before being cut off on that request. Second, the dynamic nature of social media data presented challenges - some videos included in our initial network analysis were later unavailable for content coding due to accounts being banned or made private, a particular issue in communities discussing sensitive topics like substance use. Third, methodological challenges arose around identifying substance-related content, as users often employ common words as coded references to substances, making it difficult to definitively categorize some content as substance-related versus general discussion. This challenge was compounded by the platform’s rapid evolution of slang and terminology. Additionally, some videos contained duplicate hashtags, which could affect network metrics, and less popular hashtags may be underrepresented in our sample despite efforts to ensure broad coverage. These limitations highlight the challenges of studying substance-related content on social media platforms and suggest the need for complementary research approaches to fully understand these communities.
Appendix A Appendix
Appendix B Full Hashtag Network with Communities
Figure 6 shows a visualization of the complete hashtag co-occurrence network, containing 2,333 nodes (hashtags) and 46,990 edges. The nodes are colored according to their community membership as determined by our semi-automated community detection process. The network layout was generated using a Fruchterman-Reingold algorithm, which positions nodes with stronger connections closer together. The central position of recovery-focused hashtags like “#addiction”, “#recovery”, and “#harmreduction” is clearly visible, demonstrating their role as bridges between communities. Node sizes are scaled according to their degree centrality, highlighting the most connected hashtags in the network. This visualization helps illustrate both the small-world properties of the network (clustering coefficient 0.52, average path length 2.20) and the organic formation of thematic communities around substance-related discourse on TikTok.
Figure 6: The overall network of 2,333 hashtags related to substance use with colors representing each community as shown in the legend
Appendix C Network Centrality Measures per Community
Table C presents detailed network statistics and centrality measures for both the complete network and each of the 17 identified communities including Misc. For each community, we report basic network metrics (number of nodes, edges, average degree, weighted degree, clustering coefficient, and average shortest path length) as well as the top 5 most central nodes according to four different centrality measures (betweenness, closeness, degree, and eigenvector centrality). The whole network shows high connectivity (average degree 40.28) and clustering (0.52), with core hashtags like “addiction”, “recovery”, and “sober” consistently ranking highly across all centrality measures.
Communities
Documentation of
Use for
Social Media
Informational Content Other Promotional Content Recovery Advocacy
Satirical and
Relatable Content
Trip Reports Total
Platform 8 (4.0%) 20 (9.9%) 25 (12.4%) 15 (7.4%) 59 (29.2%) 74 (36.6%) 1 (0.5%) 202
Health Conditions 0 (0.0%) 20 (11.4%) 17 (9.7%) 3 (1.7%) 100 (56.8%) 34 (19.3%) 2 (1.1%) 176
Awareness and Advocacy 1 (0.6%) 29 (16.8%) 12 (6.9%) 1 (0.6%) 99 (57.2%) 29 (16.8%) 2 (1.2%) 173
Misc 4 (3.2%) 8 (6.4%) 24 (19.2%) 7 (5.6%) 42 (33.6%) 39 (31.2%) 1 (0.8%) 125
Commonly Misused Substances 1 (0.9%) 13 (11.8%) 20 (18.2%) 1 (0.9%) 39 (35.5%) 36 (32.7%) 0 (0.0%) 110
Consumption Method 3 (3.0%) 11 (11.1%) 12 (12.1%) 14 (14.1%) 22 (22.2%) 37 (37.4%) 0 (0.0%) 99
Emotions and Feelings 4 (4.1%) 8 (8.2%) 14 (14.3%) 4 (4.1%) 32 (32.7%) 36 (36.7%) 0 (0.0%) 98
Identity and Community 2 (2.1%) 8 (8.5%) 9 (9.6%) 2 (2.1%) 51 (54.3%) 20 (21.3%) 2 (2.1%) 94
Humor 0 (0.0%) 2 (2.6%) 4 (5.3%) 1 (1.3%) 8 (10.5%) 60 (78.9%) 1 (1.3%) 76
Other Substances 2 (3.1%) 16 (24.6%) 11 (16.9%) 0 (0.0%) 21 (32.3%) 15 (23.1%) 0 (0.0%) 65
Substance Effects 7 (10.8%) 4 (6.2%) 13 (20.0%) 4 (6.2%) 6 (9.2%) 28 (43.1%) 3 (4.6%) 65
Location 3 (5.0%) 4 (6.7%) 13 (21.7%) 2 (3.3%) 9 (15.0%) 28 (46.7%) 1 (1.7%) 60
Alcohol 8 (13.6%) 6 (10.2%) 6 (10.2%) 0 (0.0%) 19 (32.2%) 18 (30.5%) 2 (3.4%) 59
Tobacco/Nicotine 2 (3.4%) 9 (15.3%) 5 (8.5%) 12 (20.3%) 16 (27.1%) 15 (25.4%) 0 (0.0%) 59
Occupation 1 (1.9%) 10 (19.2%) 9 (17.3%) 1 (1.9%) 22 (42.3%) 9 (17.3%) 0 (0.0%) 52
Cannabis 0 (0.0%) 3 (8.1%) 8 (21.6%) 0 (0.0%) 4 (10.8%) 21 (56.8%) 1 (2.7%) 37
Cognitive Enhancement 1 (4.0%) 7 (28.0%) 3 (12.0%) 7 (28.0%) 4 (16.0%) 3 (12.0%) 0 (0.0%) 25
Table 4: Distribution of communities corresponding with each hashtag used in the 351 analyzed videos per content topic.
Network #N #E Avg D Avg WD Avg C Avg SP Betweenness Closeness Degree Eigenvector
Whole Network 2333 46990 40.28 513.84 0.52 2.20 addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction
Awareness & Advocacy 279 2164 15.51 334.30 0.89 1.98 harmreduction, wedorecover, harmreductionsaveslives, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, narcansaveslives
Health Conditions 277 1522 10.99 224.09 0.85 2.02 addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, detox
Emotions & Feelings 172 305 3.55 20.65 0.31 2.39 peace, onedayatatime, friend, dream, serenity peace, onedayatatime, dream, serenity, friend peace, onedayatatime, dream, friend, serenity peace, onedayatatime, dream, serenity, friend
Platform 183 146 1.60 39.58 0.12 2.44 pingtok, dr00gtiktok, pingertok, chat, fyp pingtok, dr00gtiktok, pingertok, fyp, fyp:) pingtok, dr00gtiktok, pingertok, chat, stonertokfyp pingtok, dr00gtiktok, pingertok, chat, fyp
Identity & Community 155 97 1.25 8.22 0.09 3.21 cleanandsober, na, pwud, 30s, women pwud, na, cleanandsober, women, lgbt cleanandsober, pwud, na, 30s, mujeres cleanandsober, pwud, na, lgbtqia, calisober
Commonly Misused Substances 152 553 7.28 60.47 0.38 2.47 opioids, opiates, coke, molly, fent opioids, coke, opiates, molly, uppers opioids, opiates, coke, molly, fent opioids, coke, opiates, fent, molly
Location 116 19 0.33 1.26 0.00 1.80 chinatown, paradise, yucatan, arnolds, filmteyvatislands chinatown, paradise, india, china, texas chinatown, paradise, yucatan, arnolds, india chinatown, india, china, texas, egypt
Cannabis 51 229 8.98 39.25 0.55 2.04 maryjane, weed, cannabis, skunk, herb maryjane, weed, cannabis, herb, grass maryjane, weed, cannabis, herb, grass maryjane, cannabis, weed, herb, grass
Alcohol 43 92 4.28 56.19 0.41 1.79 drunk, tipsy, hammered, shitfaced, bars drunk, hammered, tipsy, shitfaced, bars drunk, hammered, tipsy, shitfaced, bars drunk, hammered, tipsy, bars, shitfaced
Cognitive Enhancement 24 38 3.17 45.00 0.46 1.64 nootropics, modafinil, smartdrugs, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus
Substance Effects 41 33 1.61 23.76 0.16 2.42 high, wasted, energy, blackout, clarity high, wasted, lit, crash, shadowpeople high, wasted, blackout, lit, downers high, wasted, lit, downers, crash
Consumption Method 41 62 3.02 89.80 0.30 1.86 smoke, vape, smoking, vaper, vaping smoke, vaper, vape, vaping, vaped smoke, vaper, vape, vaping, vaped smoke, vaper, vaping, vaped, vape
Tobacco Nicotine 31 115 7.42 68.90 0.70 1.70 ecigs, ecig, cigarette, nicotine, juul ecigs, cigarette, ecig, nicotine, juul ecigs, cigarette, ecig, nicotine, juul cigarette, ecig, ecigs, nicotine, juul
[…]
Occupation 73 5 0.14 0.93 0.00 1.50 copilots, truckdrivers, actress, nursing, outreachworker copilots, truckdrivers, aviation, construction, creator copilots, truckdrivers, aviation, work, construction copilots, aviation, construction, creator, truckdrivers
Misc 565 2367 8.38 139.01 0.29 2.46 cat, fire, rock, g, pink cat, fire, g, rock, pink cat, fire, g, rock, x cat, fire, g, x, rock
Table 5: Centrality measures for the whole network and each community. #N=Number of Nodes, #E=Number of Edges, Avg D=Average Degree, Avg WD=Average Weighted Degree, Avg C=Average Clustering, Avg SP=Average Shortest Path. The top five nodes are presented per centrality measure.
Appendix D Hashtag Community Distribution per Content Topic
Table 4 presents a cross-tabulation of hashtag communities and video content topics, allowing us to examine how different communities align with various types of content. This analysis reveals notable patterns in how hashtags from different communities are employed across content types. Recovery Advocacy content, the most prevalent in our dataset, shows strong representation across multiple hashtag communities, particularly drawing from the Awareness & Advocacy and Health Conditions communities. Satirical content, our second most common category, shows interesting patterns of hashtag usage, frequently incorporating hashtags from Platform and Humor communities while also drawing from more serious communities - suggesting complex layering of messaging in humorous content. Informational content shows strong alignment with Professional and Health Condition hashtags, while Promotional content primarily uses hashtags from Consumption Method and Tobacco/Nicotine communities. This distribution helps validate our community detection approach by demonstrating meaningful alignment between hashtag communities and content types, while also revealing how creators strategically combine hashtags from multiple communities to reach diverse audiences.

### v2

@@
 From #Dr00gtiktok to #harmreduction: Exploring Substance Use Hashtags on TikTok
-Abstract
-The rise of TikTok as a primary source of information for youth, combined with its unique short-form video format, creates urgent questions about how substance use content manifests and spreads on the platform.
-This paper provides the first in-depth exploration of substance use-related content on TikTok, covering all major substance categories as classified by the Drug Enforcement Agency.
-Through social network analysis and qualitative coding, we examined more than 2,333 hashtags across 39,509 videos, identified 16 distinct hashtag communities and analyzed their interconnections and thematic content.
-Our analysis revealed a highly interconnected small-world network where recovery-focused hashtags like “#addiction”, “#recovery”, and “#sober” serve as central bridges between communities.
-Through manual coding of 351 representative videos, we found that Recovery Advocacy content (33.9%) and Satirical content (28.2%) dominate, while direct substance depiction appears in only 26% of videos, with active use shown in just 6.5% of them.
-This suggests TikTok functions primarily as a recovery support platform rather than a space promoting substance use.
-We found strong alignment between hashtag communities and video content, indicating organic community formation rather than attempts to evade content moderation.
-Our findings inform how platforms can balance content moderation with preserving valuable recovery support communities, while also providing insights for the design of social media-based recovery interventions.
-Introduction
-Social media platforms have become the primary lens through which young people view and understand the world, with TikTok emerging as a dominant force among youth - reaching 63% of teens, with 17% reporting near-constant use (Vogels, Gelles-Watnick, and Anderson 2023).
-More concerning is that TikTok has become a primary news source for this demographic, often replacing Google for information-seeking (Walker and Matsa 2023).
-Figure 1: Examples of TikTok videos that can be found in the substance use communities based on hashtag use
-This shift in information consumption patterns takes on particular urgency when considering content related to illicit substances.
-While overall drug use among 8th, 10th, and 12th graders showed promising declines during COVID-19 social distancing (National Institute on Drug Abuse 2022), this drop in usage may have masked how drug-related discussions were shifting online.
-Social media platforms can rapidly disseminate trends and behaviors through psychological processes like social comparison and fear of missing out (FOMO) (Montag, Yang, and Elhai 2021), with adolescents often imitating influencers and peers they follow.
-Although most social media platforms use content moderation, their rapid and viral nature can make it challenging to effectively regulate exposure to harmful content (Boyd 2014; Lookingbill and Le 2024), potentially amplifying dangerous trends.
-Therefore, understanding such emerging trends on social media platforms such as TikTok, Instagram, and Reddit (Rutherford et al.
-2022a, 2023; Tan and Weinreich 2021; Bouzoubaa, Young, and Rezapour 2024) becomes more critical.
-For instance, the “xandemic” between 2016-2018, characterized by high school students documenting the recreational use of Xanax to the point of blackout, exemplifies how social media trends can blur the lines between casual and harmful behaviors and even normalize dangerous substance use behaviors (Fuller et al.
-2024).
-Similarly, platform-specific hashtag “#pingtok” - a portmanteau of the slang term for an MDMA pill, pinger, and TikTok - which has accumulated over 26,000 posts at the time of writing, demonstrates how substance-related content can become embedded in platform culture, with its own specific terminology and community.
-These trends are especially concerning given robust evidence that early exposure to substance use significantly increases the risk of developing substance use disorders (SUD) in adulthood (McCabe et al.
-2022; Volkow and Wargo 2022; Odgers et al.
-2008).
-Previous research on substance-related content on TikTok has typically focused on either specific substances such as cannabis or nicotine (Rutherford et al.
-2023, 2022a; Morales, Fahrion, and Watkins 2022; Tan and Weinreich 2021) or isolated trending phenomenon like “#pingtok” (Whelan, Noller, and Ward 2024).
-While some researchers have examined hashtag networks and content moderation strategies around other sensitive topics like eating disorders (Bickham et al.
-2024) and how hashtags create communities around shared experiences (Eriksson Krutrök 2021), there remains limited understanding of how hashtags shape substance-related discourse on TikTok.
-Moreover, no studies have comprehensively examined how substance-related content manifests across TikTok’s ecosystem of communities, particularly through its short-form video format and hashtag-driven discovery system (Abidin 2020).
-To address this gap, we draw from the Drug Enforcement Agency’s comprehensive drug classification system and examine content from every major substance category using a mixed-methods investigation of TikTok content.
+Abstract.
+TikTok has emerged as a major source of information and social interaction for youth, raising urgent questions about how substance use discourse manifests and circulates on the platform.
+This paper presents the first comprehensive analysis of publicly visible, algorithmically surfaced substance-related content on TikTok, drawing on hashtags spanning all major substance categories.
+Using a mixed-methods approach that combines social network analysis with qualitative content coding, we examined 2,333 substance-related hashtags, identifying 16 distinct hashtag communities and characterizing their structural and thematic relationships.
+Our network analysis reveals a highly interconnected small-world structure in which recovery-focused hashtags such as #addiction, #recovery, and #sober serve as central bridges between communities.
+Qualitative analysis of 351 representative videos shows that Recovery Advocacy content (33.9%) and Satirical content (28.2%) dominate, while direct substance depiction appears in only 26% of videos, with active use shown in just 6.5% of them.
+These findings suggest that the algorithmically surfaced layer of substance-related discourse on TikTok is predominantly oriented toward recovery, support, and coping rather than explicit promotion of substance use.
+We further show that hashtag communities and video content are closely aligned, indicating that substance-related discourse on TikTok is shaped through organic community formation within platform affordances rather than widespread adversarial evasion of moderation.
+This work contributes to social computing research by showing how algorithmic visibility on TikTok shapes the organization of substance-related discourse and the formation of recovery and support communities.
+TikTok, Substance Use, Social Media, Hashtags, Generative AI, Socio-technical System, Governance
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/25/06
+1. Introduction
+Social media platforms have become critical and contested spaces for health discourse, particularly around stigmatized topics like substance use (russell_using_2021; Johnson et al., 2025).
+TikTok’s algorithmic recommendation system has demonstrated serious harms: amplifying health misinformation (Yeung et al., 2022), promoting eating disorders (Blackburn and Hogg, 2024), and contributing to mental health crises among young users (Conte et al., 2025).
+Yet these same algorithmic affordances enable the formation of peer support communities around substance use recovery, harm reduction, and lived experience sharing (russell_using_2021; Goyal et al., 2025).
+This tension reflects broader dynamics of platform governance, in which algorithmic visibility simultaneously mediates risk, care, and community formation in stigmatized health contexts, yet remains underexplored for substance use discourse, where both risks and potential benefits are acute.
+Figure 1. Examples of TikTok videos illustrating distinct substance-related content identified through hashtags.
+Understanding how substance use content manifests on platforms like TikTok is critical for multiple stakeholders.
+Early exposure to substance use significantly increases the risk of developing substance use disorders (SUD) (McCabe et al., 2022; Volkow and Wargo, 2022; Odgers et al., 2008), making youth exposure a public health concern.
+However, platforms also serve people already living with SUD who may lack offline support systems (russell_using_2021; Bouzoubaa et al., 2024b; Goyal et al., 2025).
+Previous trends like the “xandemic” (2016-2018), characterized by high school students documenting the recreational use of Xanax to the point of blackout, and #pingtok, a portmanteau of the slang term for an MDMA pill, pinger, and TikTok, demonstrate how substance-related content can become embedded in platform culture (Fuller et al., 2024; rutherford_changes_2022).
+What remains unclear is how substance normalization and recovery-support dynamics coexist, intersect, and vary across substances and communities.
+Previous research on substance-related content on TikTok has typically focused on either specific substances such as cannabis or nicotine (rutherford_turnttrending_2023; rutherford_changes_2022; morales_nicotineaddictioncheck_2022; tanPuffBarHowTop2021) or isolated trending phenomena such as #pingtok (Whelan et al., 2024).
+While some researchers have examined hashtag networks and content moderation strategies around other sensitive topics like eating disorders (Bickham2024HiddenInPlainSight) and how hashtags create communities around shared experiences (eriksson_krutrok_algorithmic_2021), there is limited comprehensive knowledge of how substance-related discourse is organized across TikTok’s hashtag-driven ecosystem.
+In contrast to prior work that examines individual substances, isolated trends, or moderation outcomes, we take a platform-wide, cross-substance perspective that captures how substance discourse is structurally organized and socially enacted across TikTok’s hashtag ecosystem.
+To address this gap, we draw on an established drug classification system (United States Drug Enforcement Administration, 2025) in order to examine content from every major substance category on TikTok.
 Our research questions are:
@@
   • RQ2: What are the topics and key characteristics of the videos associated with the hashtags in these communities?
-  • RQ3: What are the effects of substance visibility on topic distribution and engagement metrics?
-For this purpose, we developed a list of 585 seed hashtags related to substance use and collected more than a thousand videos for each of them, resulting in a dataset of 504,294 hashtags and 272,910 videos  [^1] Datasets, including seed hashtags, and code are available at https://github.com/social-nlp-lab/explore˙tiktok˙su.
-Videos are not directly provided but can be accessed via URLs..
-After filtering relevant hashtags, we identified 2,333 exemplary hashtags covering the broad categories of substance use, and 351 videos for qualitative coding.
-Using a combination of network analysis of hashtag co-occurrence and in-depth content analysis, our analysis showed a highly interconnected small-world network where recovery-focused hashtags like “#addiction”, “#recovery”, and “#sober” serve as central bridges between communities.
-Within these communities, we found that Recovery Advocacy content (33.9%) and Satirical content (28.2%) dominate, while direct substance depiction appeared in only 26% of videos, with active use shown in just 6.5%.
-This suggests TikTok functions primarily as a recovery support platform rather than a space promoting substance use.
-Through this work, we contribute to social computing and substance use literature by providing the first comprehensive mapping of substance-related content on TikTok.
-Our findings can inform platform policies around content moderation strategies that preserve supportive recovery communities while better identifying potentially harmful content, guide the development of social media-based recovery support interventions, and help design youth substance use prevention efforts that account for how substance-related content actually manifests on TikTok.
-Related Works
+  • RQ3: How does substance visibility, operationalized through engagement and algorithmic surfacing, affect topic distribution and engagement metrics?
+We address these questions through a mixed-methods analysis combining social network analysis of hashtag co-occurrence patterns with qualitative content coding.
+Analyzing 2,333 exemplar hashtags and coding 351 representative videos, we examine both the structural organization of substance-related discourse and the actual content users create.
+Our network analysis reveals a highly interconnected discourse where recovery-focused hashtags like #addiction, #recovery, and #harmreduction serve as bridges between thematic communities.
+Qualitative coding shows that Recovery Advocacy (33.9%) and Humor-based coping (28.2%) make up the majority of videos, while direct substance depiction appears in only 26% of videos, with active use shown in just 6.5%.
+Our findings suggest that substance-related content on TikTok, within the publicly visible, algorithmically surfaced layer accessible to observation, is predominantly oriented toward support rather than promotion or glamorization.
+Through this work, we contribute to social computing and substance use literature by providing the first comprehensive mapping of substance-related content on TikTok across all major drug categories, revealing 16 distinct hashtag communities within a highly interconnected small-world network.
+Second, we demonstrate that hashtag communities exhibit organic thematic coherence rather than adversarial evasion of content moderation, suggesting these support communities form within rather than against platform affordances.
+Third, we show that recovery advocacy (33.9%), humor-based coping (28.2%), and educational information (13.1%) collectively make up most of the content, with active substance use depiction making up only 6.5% in algorithmically surfaced public discourse.
+Together, these contributions advance social computing research on algorithmic visibility, health discourse, and community formation under platform governance constraints.
+Finally, we derive design implications for content moderation strategies that must balance harm prevention with preservation of beneficial peer support communities.
+These findings inform platform governance approaches, the development of social media-based recovery interventions, and youth substance use prevention efforts that account for how substance-related content actually manifests on TikTok.
+2. Related Work
 As digital platforms become ubiquitous, understanding their impact, especially on young people’s health behaviors regarding substance use, is more crucial than ever.
-On the one hand, social media emerged as a medium for facilitating sensitive discussions and disclosure of lived experiences, including mental health and eating disorders (Rutherford et al.
-2022a; Bickham et al.
-2024; Bouzoubaa et al.
-2024).
+On the one hand, social media emerged as a medium for facilitating sensitive discussions and disclosure of lived experiences, including mental health and eating disorders (rutherford_changes_2022; Bickham2024HiddenInPlainSight; Bouzoubaa et al., 2024a).
 On the other hand, several studies have examined the relationship between online peer groups, social media homophily, and substance use, finding positive correlations between substance use and social media engagement.
-For example, Willoughby et al.
-(2024) found that young adults reported an increased intention to use cannabis after exposure to pro-cannabis messages online.
-Similarly, Sun et al.
-(2023) showed that exposure to e-cigarette advertisements on social media platforms is associated with adolescents’ e-cigarette usage in later years.
-Miller et al.
-(2021) examined how online and offline peer interactions influence substance use behaviors and found that online belonging and social media homophily predicted regular stimulant and opioid use among substance users.
-Hanson et al.
-(2013) have studied online discussions related to stimulant drug use on Twitter and found a high correlation between the popularity of such topics and final exam periods.
-Geusens et al.
-(2023) found that higher levels of alcohol and marijuana consumption predicted later social media posts about those substances, but posting did not predict later substance use.
-Sun et al.
-(2023) found a positive correlation between early exposure to e-cigarette advertisement and subsequent lifetime usage.
-Hashtags, a medium for interactions on platforms like Facebook or Twitter, have been studied comprehensively in the past (Mousavi and Ouyang 2021; Ince, Rojas, and Davis 2017; Small 2011; Rezapour et al.
-2017).
-People can contribute to the narrative by hash-tagging their own personal experiences and thoughts (Yang 2016).
-At their core, hashtags facilitate content discovery and community creation by allowing users to contribute to the discussion through keyword tagging (Small 2011).
+For example, willoughby_exposure_2024 found that young adults reported an increased intention to use cannabis after exposure to pro-cannabis messages online.
+Similarly, sun_longitudinal_2023 showed that exposure to e-cigarette advertisements on social media platforms is associated with adolescents’ e-cigarette usage in later years.
+miller_online_2021 examined how online and offline peer interactions influence substance use behaviors and found that online belonging and social media homophily predicted regular stimulant and opioid use among substance users.
+hanson_tweaking_2013 have studied online discussions related to stimulant drug use on Twitter and found a high correlation between the popularity of such topics and final exam periods.
+geusens_first_2023 found that higher levels of alcohol and marijuana consumption predicted later social media posts about those substances, but posting did not predict later substance use.
+sun_longitudinal_2023 found a positive correlation between early exposure to e-cigarette advertisement and subsequent lifetime usage.
+Hashtags, a medium for interactions on platforms like Facebook or Twitter, have been studied comprehensively in the past (mousavi_detecting_2021; ince_social_2017; small_what_2011; Rezapour et al., 2017).
+At their core, hashtags facilitate content discovery and community formation by enabling users to collectively construct and navigate shared narratives through keyword tagging (Yang, 2016; small_what_2011).
 Previous studies have examined substance use discourse across various social media platforms, contributing to our understanding of how substance use behaviors manifest online.
-For instance, on Instagram, investigating “#studydrugs” demonstrated a predominantly positive portrayal of stimulant use for academic performance enhancement (Petersen et al.
-2021).
-On TikTok, content creators deliberately employ hashtags like “#fyp” (For You Page) to increase visibility and maximize audience reach within TikTok’s recommendation system (Abidin 2020).
-The platform’s ranking algorithms connect users based on hashtag engagement patterns, creating what Eriksson Krutrök (2021) calls “algorithmic closeness” - subcommunities united through shared content interests.
+For instance, on Instagram, investigating #studydrugs demonstrated a predominantly positive portrayal of stimulant use for academic performance enhancement (petersen_studydrugspersuasive_2021).
+On TikTok, content creators deliberately employ hashtags like #fyp (For You Page) to increase visibility and maximize audience reach within TikTok’s recommendation system (abidin_mapping_2020).
+The platform’s ranking algorithms connect users based on hashtag engagement patterns, creating what eriksson_krutrok_algorithmic_2021 calls “algorithmic closeness” - subcommunities united through shared content interests.
 This algorithmic dynamic shapes how substance-related discourse spreads and evolves on TikTok.
-For example, research examining TikTok hashtags like “#puffbar” has revealed concerning patterns of youth-oriented substance marketing and normalization (Tan and Weinreich 2021), while analysis of “#stonertok” and “#420vibes” showed the majority of cannabis-related videos positively portray substances (Rutherford et al.
-2022b).
+For example, research examining TikTok hashtags like #puffbar has revealed concerning patterns of youth-oriented substance marketing and normalization (tanPuffBarHowTop2021), while analysis of #stonertok and #420vibes showed the majority of cannabis-related videos positively portray substances (rutherford_getting_2022).
 Our study advances existing research by employing a network-based and qualitative approach to explore the interactions and communities associated with substance-related hashtags on TikTok.
-While previous research has thoroughly documented individual communities, the dynamics of how these groups interact across the platform’s broader ecosystem remain understudied.
-Our work can help social media platforms like TikTok refine their policies to better distinguish between harmful content requiring removal and harm reduction information that should be preserved, contributing to more effective content moderation.
-Methods
-Data
-Seed Hashtag Selection: To develop an initial set of seed hashtags related to substance use for the data collection, we first identified hashtags previously used in studies of substance-related content on social media platforms (Petersen et al.
-2021; Russell et al.
-2022; Rutherford et al.
-2023, 2022b; Singh and Wu 2021) as well as searching for hashtags on the platform associated with substance names (both formal and slang) listed on the Drug Enforcement Agency’s Fact Sheet (United States Drug Enforcement Administration 2025).
+While prior research has examined individual substances or discrete hashtag communities, the structural relationships and interactions among these communities within TikTok’s broader ecosystem remain underexplored.
+Rather than focusing on message content or sentiment alone, we examine how substance-related discourse is structurally organized and interconnected at the platform level.
+Our work provides empirical grounding for understanding algorithmic visibility, community formation, and the coexistence of harm and support within substance-related discourse, informing not only content moderation but also platform governance, intervention design, and the development of evidence-based recommendations for public health and recovery-oriented stakeholders.
+3. Methods
+Figure 2. Data selection and pre-processing pipeline to create the filtered set.
+3.1. Data
+3.1.1. Seed Hashtag Selection:
+To develop an initial set of seed hashtags related to substance use for the data collection, we first identified hashtags previously used in studies of substance-related content on social media platforms (petersen_studydrugspersuasive_2021; russell_analyzing_2022; rutherford_turnttrending_2023; rutherford_getting_2022; singh_sentiment_2021) as well as hashtags associated with substance names (including both formal and slang terms) listed on the Drug Enforcement Agency’s Fact Sheet (United States Drug Enforcement Administration, 2025).
 Using this set, we conducted iterative searches on TikTok to identify additional relevant hashtags by examining co-occurring hashtags within the content.
-This process continued until we reached saturation, resulting in a final set of 585 substance-use-related hashtags.
+This process continued until we reached saturation (no substantively new substance-related hashtags yielded), resulting in a final set of 585 substance-use-related hashtags.
 [^1] The list will be shared upon the paper’s acceptance.
-Data Collection: Using the Unofficial TikTok API [^3] https://github.com/davidteather/TikTok-Api, we collected data for around 1,000 algorithmically served videos for each of the 585 seed hashtags.
-This resulted in a dataset of 275,291 videos and a total of 504,294 unique hashtags.
-Figure 2: Data selection and pre-processing pipeline to create the filtered set.
-Data Preprocessing and Network Refinement: To enhance the feasibility of our analysis and reduce noise from potentially irrelevant content, we filtered the data to include only videos that featured at least two of the seed hashtags.
-This filtering step resulted in a refined dataset of 39,602 videos and 99,727 unique hashtags (Figure 2).
-To further clean the data, we constructed a co-occurrence network where nodes represent individual hashtags and edges represent their co-occurrence in videos.
-More specifically, edge weights were defined by the frequency with which two hashtags appeared together in videos.
-We then employed a two-stage filtering process to focus on the most relevant and interconnected hashtags.
-First, we removed hashtags with low connectivity by filtering out nodes in the bottom 25th percentile of degree centrality.
-This step eliminated peripheral hashtags with minimal connections to the broader substance use discourse, removing those that were incidentally rather than substantively related to substance use discussions.
-Next, we applied k-core decomposition ($k=10$) to identify the densely connected core of the network (Seidman and Foster 1978).
+3.1.2. Video Collection:
+Using the Unofficial TikTok API [^2] https://github.com/davidteather/TikTok-Api, we collected up to 1,000 publicly available, algorithmically surfaced videos for each of the 585 seed hashtags during the data collection period.
+This resulted in an initial, raw dataset of 272,910 videos.
+These videos contained a total of 504,294 unique hashtags (including both substance-related terms and general hashtags that co-occurred with them).
+3.1.3. Network Construction and Refinement:
+To enhance the feasibility of our analysis and reduce noise from potentially irrelevant content, we filtered the data to include only videos that featured at least two of the seed hashtags, ensuring that retained videos reflected explicit engagement with substance-related discourse rather than incidental hashtag use.
+This filtering step resulted in a refined dataset of 99,727 unique hashtags (Figure 2).
+To further clean the data, we constructed a hashtag co-occurrence network where nodes represent individual hashtags and edges represent their co-occurrence in the same video.
+Edge weights were defined as the frequency with which two hashtags appeared together in videos.
+We then employed a two-stage network refinement process to focus on the most relevant and interconnected hashtags.
+First, we removed hashtags with low connectivity by filtering out nodes in the bottom 25th percentile of degree centrality (borgatti2005centrality).
+This step eliminated peripheral hashtags with minimal connections to the broader substance use discourse, removing those that were weakly connected or contextually incidental rather than substantively related to substance use discussions.
+Next, we applied k-core decomposition ($k=10$) to identify the densely connected core of the network (Seidman and Foster, 1978).
 K-core decomposition recursively removes nodes with degrees less than $k$ until all remaining nodes have at least $k$ connections, ensuring a cohesive network structure.
-This threshold was chosen based on exploratory analysis, which revealed that $k=10$ produced the most structurally cohesive subset of hashtags while retaining a sufficient number of nodes and edges to analyze the discourse.
+The value of $k=10$ was selected based on exploratory analysis, which indicated that this threshold preserved network connectivity while excluding sparsely connected hashtags.
 The resulting refined network of hashtags contained 2,333 nodes and 46,990 edges, representing the most interconnected hashtags in substance use-related content.
@@
 Misc Any tag that does not fit into the above categories. #cookies, #foundersday
-Table 1: Categories for semantically classifying hashtags from TikTok content, their definitions, and example hashtags.
-Substance Use Hashtags Network
-Community Detection: We first applied established community detection algorithms including Clauset-Newman-Moore (Clauset, Newman, and Moore 2004) and Louvain method (Blondel et al.
-2008), but these approaches did not yield clusters that meaningfully captured substance-related discussions.
+Table 1. Categories for semantically classifying hashtags from TikTok content, their definitions, and example hashtags.
+3.2. Substance Use Hashtags Network Analysis (RQ1)
+3.2.1. Community Classification:
+We first applied established community detection algorithms, including Clauset-Newman-Moore (Clauset et al., 2004) and Louvain method (Blondel et al., 2008), but these approaches did not result in clusters that meaningfully captured substance-related discussions, as the resulting communities were driven primarily by network density rather than thematic or semantic coherence.
 Therefore, we used a hybrid approach combining manual and automated methods to identify thematic groups within the hashtags.
-We first manually annotated a sample of 251 hashtags and iteratively established thematic categories.
+We first manually annotated a sample of 251 hashtags drawn from prior work, manual TikTok searches and DEA fact sheets (United States Drug Enforcement Administration, 2025) and iteratively established substance-related content categories.
+This iterative process was informed by prior work on substance use discourse and refined through repeated examination of hashtag usage and contexts.
 As a result, we identified 16 categories representing topics related to substance use and consumption, identity, and community, health and emotion as well as awareness and advocacy.
@@
 The final set of categories and their definitions are shown in Table 1.
-These categories were then used to prompt GPT-4o (OpenAI 2024) to classify hashtags.
-When evaluated against human-labeled data, our best-performing prompt achieved an accuracy of $89.36\%$, demonstrating strong reliability in the thematic categorization.
+These categories were then used to prompt GPT-4o (OpenAI, 2024) to classify hashtags.
+When evaluated against human-labeled data, our best-performing prompt achieved an overall accuracy of $89.36\%$, demonstrating strong reliability in the thematic categorization.
 We then used this prompt to classify the full set of 2,333 hashtags.
 The resulting labeled hashtags were mapped back onto the network, creating themed communities that reflect different aspects of substance use discourse on TikTok.
-Network Analysis: For both the overall network and each community, we calculated standard network centrality measures including betweenness (Freeman 1977), closeness (Sabidussi 1966), degree (Freeman 1978), and eigenvector centralities (Bonacich 1972).
-These measures help identify influential hashtags and understand the structure of discourse within and across themes.
-Additionally, we computed network-level metrics such as average degree, clustering coefficient, and average shortest path length to characterize the overall connectivity patterns and community structure of substance use-related discourse on TikTok.
-Characterizing Video Content
-Codebook Development: To establish the codebook for coding substance-related TikTok videos, two researchers independently open-coded 160 randomly selected videos.
-Through this initial coding process, researchers identified nine distinct video topics and coded for substance presence and usage.
-The physical presence of drugs and related substances was of interest to us in determining which communities and topics were more engaged in depicting substances.
-The intercoder reliability was strong for coding substance visibility ($\kappa=0.87$) but only fair for video topics ($\kappa=0.35$).
-After discussing initial disagreements, the topics were refined from nine to seven distinct categories.
-The coders then had an iterative process for discussing the codes with all researchers on the team to identify coherent topics of videos within the TikTok dataset, resolve any disagreements, and finalize the codebook.
-The final codebook captured two key dimensions: video topic and substance presence, as shown in Table 3.
-Video Selection: We started by selecting all videos that contained at least one of the 2,333 core hashtags, resulting in a dataset of 39,509 videos.
-To refine this dataset for detailed coding, we identified the top five hashtags per community based on their degree centrality.For each of the 80 selected hashtags, we then filtered videos using a stepwise approach:
-  1. 1.
-    50 initial videos were acquired from our dataset for each of the 80 hashtags.
-Ten of each were acquired from each quintile of viewership for the hashtag, achieving representation regardless of virality.
-  2. 2.
-    Each video was evaluated against three key criteria:
-    • Availability: The video is accessible and not moderated or removed by TikTok.
-    • Language: The video features English as the primary spoken or written language.
-    • Substance Relevance: The video clearly depicts substance-related content, whether through visual cues, discussion, or behaviors.
-We continued this process until we identified five videos per hashtag, one per each quintile of engagement, that met all three criteria.
-In total 977 links were coded, with 852 videos available (87.2%), of those 695 videos were in English (81.5%), and of those 351 videos were substance-related (50.5%).
-We conducted an inductive content analysis (Elo and Kyngäs 2008) on these 351 selected videos based on the codebook to ensure our final dataset was representative of the substance-use discourse across different communities (See Table 2).
-Findings
+3.2.2. Network Analysis:
+For both the overall network and each community, we calculated standard network centrality measures including betweenness (Freeman, 1977), closeness (Sabidussi, 1966), degree (Freeman, 1978), and eigenvector centralities (Bonacich, 1972).
+These widely used measures capture complementary aspects of structural prominence, including reachability, connectivity, and association with highly connected hashtags and help identify influential hashtags in terms of their structural roles within the discourse network.
+Additionally, we computed network-level metrics such as average degree, clustering coefficient, and average shortest path length to characterize overall connectivity, cohesion, and small-world properties of substance use–related discourse on TikTok, both at the global network level and within individual communities.
+3.3. Qualitative Analysis of Video Content (RQ2)
+Building on the hashtag communities identified in RQ1, we analyzed video content associated with the 2,333 core hashtags organized into 16 thematic clusters (excluding the Misc cluster).
+While network analysis reveals how substance-related discourse is structurally organized, it cannot capture the nature of content within these communities (e.g., whether videos promote use, support recovery, or serve other functions).
+To address RQ2, we qualitatively examined how substance-related discourse manifests within and across these communities along two dimensions: video topic and substance presence.
+3.3.1. Codebook Development
+We developed a qualitative codebook using an iterative open-coding process designed to capture salient characteristics of substance-related video content (braun2006using).
+Two members of the research team independently coded an initial set of 160 randomly selected videos drawn from the broader 39,509-video dataset.
+Through this exploratory open-coding process, the coders identified an initial set of nine video topic categories reflecting recurring themes in the content.
+In parallel, we coded for substance presence, capturing whether substances were visually depicted or actively used within videos.
+This dimension was included to assess how different communities and topics varied in their explicit depiction of substances.
+Following this initial round, the research team met to discuss discrepancies, refine category definitions, and consolidate overlapping themes.
+Through another round of independent coding, reliability was assessed on a 10% (n = 40) of the coded videos, yielding moderate agreement for video topic ($\kappa=0.51$) and substance shown ($\kappa=0.58$).
+This process resulted in a reduced set of seven video topic categories.
+Remaining disagreements were resolved through deliberative discussion among the full research team until consensus was reached.
+The 160 development videos were excluded from the final analytic sample to maintain independence between codebook development and application.
+The final codebook captured two primary dimensions: (1) video topic and (2) substance presence.
+Definitions and examples for each category are provided in Table 3.
+3.3.2. Video Sampling
+We constructed the qualitative sample through a multi-stage, stratified selection process designed to ensure coverage across communities and engagement levels.
+We started with the full set of 39,509 videos containing at least one of the 2,333 core hashtags.
+To ensure representation across the 16 hashtag communities, we first selected the top five hashtags per community based on degree centrality, yielding a total of 80 focal hashtags.
+Degree centrality was used to prioritize hashtags that were structurally prominent within each community’s discourse network.
+For each focal hashtag, we stratified associated videos into five engagement quintiles based on view count, calculated within each hashtag’s distribution.
+Our target was to code at least one video per engagement quintile per hashtag (5 videos × 80 hashtags = 400 minimum).
+To account for anticipated exclusions due to content removal, non-English language, or lack of substance relevance, we initially sampled up to 10 videos per quintile where available.
+However, many hashtags had fewer than 50 associated videos in the filtered dataset, resulting in incomplete coverage of some quintile-hashtag combinations.
+This process yielded 977 candidate videos for screening.
+Each candidate video was then evaluated against three inclusion criteria:
+  • Availability: The video was publicly accessible at the time of coding (i.e., not removed or restricted).
+  • Language: The primary spoken or written language of the video was English.
+  • Substance Relevance: The video clearly involved substance-related content, indicated through visual cues, discussion, or behaviors.
+Of the 977 candidate videos, 852 (87.2%) were publicly available at the time of analysis, 695 (81.5% of available videos) were in English, and 351 (50.5% of English videos) met the substance relevance criterion.
+These 351 videos constituted the final corpus for inductive qualitative content analysis (Elo and Kyngäs, 2008).
+The videos were divided between the two coders for independent coding.
+Coders consulted with the broader research team when encountering ambiguous cases to ensure consistent application of the codebook.
+We report descriptive frequencies for all coded dimensions.
+Descriptive statistics for the sampled videos are reported in Table 2.
+4. Findings
 Filtered Coded
@@
 Max. Date 2024-05-13 2024-05-09
-Table 2: Video statistics of the dataset filtered for RQ1 and the sub-set used for coding in RQ2.
-RQ1: Network Structure and Communities
-Analysis of hashtags network showed complex, interconnected structures with distinct communities.
-The overall network contained 2,333 nodes (hashtags) and 46,990 edges (Table C in Appendix), with an average degree of 40.28 and a weighted degree of 513.84 as listed in Table C.
-These metrics indicate substantial interconnection - each hashtag co-occurs with approximately 40 other hashtags, and when hashtags do co-occur, they appear together frequently (over 500 times on average).
-The network exhibits small-world properties characterized by a high average clustering coefficient (0.52) coupled with a short average path length (2.20) (Watts and Strogatz 1998).
-This structure suggests that while substance-related discussions form tight communities, information can still flow efficiently across the network through just two intermediate hashtags on average.
-Network centrality analyses revealed two distinct categories of influential nodes.
-The first category consists of substance-specific hashtags like “#addiction”, “#recovery”, “#sober”, “#wedorecover”, and “#harmreduction”, which scored highly across all centrality measures (betweenness, closeness, degree, and eigenvector).
-These hashtags function as critical bridges between different substance-related discussions, facilitating information flow between harm reduction, recovery, and other communities.
-The second category includes general-purpose hashtags like “#fyp” and “#mentalhealth”, which exhibit high centrality but serve primarily to connect substance-related discourse to broader TikTok conversations.
-This integration with platform-wide hashtags means substance-related content can potentially reach users not explicitly seeking such content through TikTok’s recommendation system.
-Our community detection analysis identified 16 different themes (excluding Misc, see Table 1), each revealing distinct patterns in how substance-related content is organized on TikTok (Figure 6 in Appendix).
-The largest community in this network represents Awareness and Advocacy, containing approximately 12% of all hashtags (280 nodes) and featuring highly central nodes focused on harm reduction (e.g., “#harmreduction”, “#harmreductionworks”), recovery and sobriety (e.g., “#wedorecover”, “#sober”), and public health awareness (e.g., #opioidawareness, #opioidcrisis).
-The Emotions and Feelings cluster, while smaller with 180 nodes, shows high internal connectivity (clustering coefficient of 0.61) and frequently bridges to other communities through hashtags like “#mentalhealth” and “#anxiety,” suggesting the important role of emotional support in substance-related discussions.
-The network also revealed unique patterns in how substance-specific discussions are organized.
-Communities such as Cannabis (142 nodes) and Tobacco/Nicotine (138 nodes), exhibit strong internal clustering but limited external connections.
-This isolation suggests the discussions stay within their topical boundaries.
-The Commonly-Misused Substances cluster, though relatively small (120 nodes), contains hashtags often straddling the line between awareness and potential promotion of substance use (e.g., ”#studydrugs”, “#pingertok”, “#smartdrugs”).
-The Platform-specific cluster (165 nodes) showed distinct characteristics (e.g., “#fyp”, “#viral”), with the highest average betweenness centrality, indicating its crucial role in amplifying substance-related content to broader audiences.
-Figure 3: Engagement metrics across select popular hashtags in the coded set. Note, #harmreductionsaveslives has been shorted to HSL.
-RQ2: Topics and Characteristics of Videos
-Analysis of 351 videos based on our codebook (Table 3) showed that the most prevalent type of content in the videos focused on advocacy and recovery support (33.9%, 119 videos), where creators shared personal recovery journeys, celebrated sobriety milestones, or provided support to others struggling with substance use.
-For example, one video shows a creator celebrating “100 days sober milestone and encouraging others to seek help.” Satirical or humorous content formed the second largest category (28.2%, 99 videos), with videos often using comedy to address serious topics around substance use.
-The content is typically presented in the form of acted-out skits or parodies designed to gain engagement.
-For instance, one video presented a “comedic skit about stereotypical behavior while intoxicated.”
-Informational content, typically from claimed medical or professional perspectives, comprised 13.1% (n=46) videos.
-These videos often featured “doctors explaining the effects of new medications on brain chemistry” or provided educational content about substance effects and risks.
-A smaller portion focused purely on social media engagement (3.9%, n=14 videos), with creators recording substance-related content primarily to gain followers or views.
-A sizable portion of the content is in the promotional category, appearing in 7% (n=24) videos.
-The majority of them portray a demonstration of a vaping device, Juul or lighter, and a few of them promote nootropics or study drugs.
-Notably, while discussions of substances were common, direct depiction of substances was less frequent - appearing in 25.9% (n=91) of videos, with only 23 of those videos showing active substance use.
-This suggests a tendency toward discussing rather than displaying substance use.
+Table 2. Video statistics of the dataset filtered for RQ1 and the sub-set used for coding in RQ2.
+4.1. RQ1: Network Structure and Communities
+Analysis of the hashtag network revealed a highly interconnected structure with 2,333 nodes (hashtags) and 46,990 edges (Table A.1 in Appendix), with an average degree of 40.28 and a weighted degree of 513.84.
+These metrics indicate substantial interconnection; each hashtag co-occurs with approximately 40 other hashtags, and when hashtags do co-occur, they appear together frequently (over 500 times on average).
+The network exhibits small-world properties (watts_collective_1998), characterized by a high average clustering coefficient (0.52) and a short average path length (2.20).
+This structure suggests that substance-related discussions form tightly knit clusters while remaining structurally connected through relatively short paths across the broader network.
+Network centrality analysis identified recovery-focused hashtags as structurally central.
+The five hashtags with highest betweenness centrality: #addiction (0.129), #recovery (0.098), #sober (0.058), #wedorecover (0.050), and #harmreduction (0.049), all relate to recovery and harm reduction, positioning these terms as bridges connecting otherwise disparate thematic communities.
+Community analysis identified 16 thematic clusters (excluding Misc, see Table 1), each reflecting distinct structural properties within substance-related discourse (Figure A.1 in Appendix).
+Communities varied substantially in internal cohesion.
+The largest community, Awareness and Advocacy, comprises approximately 12% of all hashtags (279 nodes) with high clustering (0.89), and includes hashtags related to harm reduction (e.g., #harmreduction, #harmreductionworks), recovery and sobriety (e.g., #wedorecover, #sober), and public health awareness (e.g., #opioidawareness, #opioidcrisis).
+Similarly, Health Conditions (277 nodes, clustering = 0.85) showed high internal cohesion around hashtags such as #addiction and #chronicpain.
+Substance-specific communities exhibit varied structural patterns.
+Cannabis (51 nodes, clustering = 0.55) and Tobacco/Nicotine (31 nodes, clustering = 0.70) show high internal clustering, while Commonly-Misused Substances (152 nodes, clustering = 0.38) and Alcohol (43 nodes, clustering = 0.41) exhibit more moderate cohesion.
+The Commonly-Misused Substances community contains hashtags that often span awareness-oriented and substance-use-oriented content (e.g., #studydrugs (petersen_studydrugspersuasive_2021), #pingertok (Whelan et al., 2024), #smartdrugs (petersen_studydrugspersuasive_2021)).
+In contrast, several communities function as cross-cutting connectors with low internal clustering.
+The Consumption Method community (41 nodes, clustering = 0.30), containing hashtags such as #smoke, #vape, and #injection, exhibit the highest average betweenness centrality among all communities, connecting discussions across substance types.
+The Platform-Specific community (183 nodes, clustering = 0.12), characterized by hashtags such as #fyp, #viral, exhibits minimal internal structure with only 3.2% of its edges occurring within the community, consistent with its function as a mechanism for algorithmic visibility rather than topical organization.
+Figure 3. Engagement metrics across select popular hashtags in the coded set. Note, #harmreductionsaveslives has been shorted to HSL.
+4.2. RQ2: Video Topics and Substance Presence
+Analysis of 351 videos using the qualitative codebook (Table 3) revealed that the most prevalent category of content focused on advocacy and recovery support (33.9%, $n=119$ videos).
+Videos in this category commonly featured creators sharing personal recovery journeys, celebrating sobriety milestones, or offering encouragement and support to others navigating substance use challenges.
+For example, one video documented a creator celebrating a “100 days sober” milestone while encouraging viewers to seek help.
+Satirical and humorous content comprised the second largest category (28.2%, $n=99$ videos).
+These videos frequently used comedy or parody, often through skits or reenactments (e.g., “comedic skit about stereotypical behavior while intoxicated”), to engage with substance-related experiences.
+Informational content, often framed from claimed medical or professional perspectives (e.g., “doctors explaining the effects of new medications on brain chemistry”), accounted for 13.1% ($n=46$) of videos and included educational explanations of substance effects, risks, or treatment options.
+A smaller subset of videos focused primarily on platform engagement (3.9%, $n=14$ videos), where substance-related themes were used mainly to attract views or followers.
+Promotional content appeared in 7.0% ($n=24$) of videos, most commonly featuring demonstrations of vaping devices, Juul or lighter, as well as occasional promotion of nootropics or so-called study drugs.
+Across all content categories, direct depiction of substances was comparatively limited.
+Substances were visually present in 25.9% ($n=91$) of videos: 19.4% ($n=68$) showed substances without active use, while only 6.6% ($n=23$) depicted active substance use.
+This pattern suggests that substance-related discourse on TikTok more frequently involves discussion, narration, or symbolic reference rather than explicit visual display.
 Dimension Category Definition N (%) Example
@@
 Other Substance-related content that doesn’t fit other categories 45 (12.8%) News clip about substance policy changes
-Table 3: Codebook for qualitative analysis of substance-related TikTok videos (N=351)
-Figure 4: Sub-networks of Awareness & Advocacy (Pink) and Health Conditions (Blue) to visualize strong overlap between hashtag usage.
-Alignment Between Hashtags and Video Content We also examined whether videos’ actual content aligns with their associated hashtag and communities.
-While each video was coded for its primary topic, we find that videos typically contain hashtags from multiple communities, revealing complex patterns in how substance-related content is tagged and disseminated.
-Advocacy, Recovery, and Support Content Videos coded primarily as advocacy and recovery content (119 videos, 33.9%), the largest thematic group in our manual coding, showed strong alignment with two major hashtag communities (see Table 4):
-  • Awareness & Advocacy (99 videos, 57.2%): Recovery-focused videos frequently used hashtags like “#harmreduction” and “#wedorecover”.
-Notably, while recovery-focused hashtags were more commonly used overall, content tagged with harm reduction messaging (particularly “#harmreductionsaveslives”) generated substantially higher user engagement across all metrics - views, likes, comments, and shares (Figure 3).
-  • Health Conditions (100 videos, 56.8%): Videos discussing recovery often included medical or health-focused tags like “#addiction”, “#chronicpain”, and “#pain” (Figure 4)
-Content of videos seen in this group are typically centered around active or past efforts to recover from substance abuse, discussed in context of cautionary tales, or implicitly imploring the viewer to regard substance abuse with regard and care due to potential harms.
-Entertainment and Platform Engagement Entertainment and humor-focused content (99 videos, 28.2%) as well as videos that documented using a substances for engagement (14 videos, 4%) similarly bridged multiple communities.
-These videos predominantly contained hashtags from the communities:
-  • Platform (82 videos, 40.6% of all coded videos): Entertainment videos heavily utilized platform-specific tags like “#fyp” and “#pingtok”
-  • Humor (60 videos, 78.9%): Videos using “#addictionhumor” and similar tags often presented substance use through comedy or satire
-  • Substance Effects (35 videos, 53.8%): Videos that were satirical or relatable often documented an effect of substance use (e.g., what it’s like being “#high” or “#stoned”).
-The subjects in these videos may not necessarily be intoxicated but could be enacting or portraying the effects.
-Content in this group still discussed topics of substance use and misuse, however the seeming salient topic was humor.
-A large proportion of videos in this group were meme-based videos, where the original content was either not drug related but brought into drug discussion with filters, captions, or additional context, or was drug related but leaned towards humor and not object lessons.
-Figure 5: Distribution of video topics across three levels of substance visibility (no substances shown, substances shown but not used, and active substance use) in TikTok videos.
-Educational and Informational Content Videos coded as informational (46 videos, 13.1%) showed connections to multiple professional and health-focused communities:
-  • Health Conditions (20 videos, 11.4%): Medical information and health impacts
-  • Awareness & Advocacy (29 videos, 16.8%): Harm reduction education
-  • Other Substances (16 videos, 24.6%) :
-  • Occupation (10 videos, 19.2%): Content from healthcare workers and other professions
-Content in this group was thematically clinical or information focused.
-Humor was used sparingly if at all.
-The vast majority of videos in this cluster used detailed captions to go over specific facts regarding drug use or were more long-form and had a health professional or recovery advocate speak at length on the topic.
-Recovery organizations were seen multiple times in the group, with high production value videos.
+Table 3. Codebook for qualitative analysis of substance-related TikTok videos (N=351)
+We next examined how videos’ content aligned with their associated hashtag and communities.
+While each video was coded for its primary topic, videos typically contained hashtags from multiple communities, revealing complex patterns in how substance-related content is tagged and disseminated.
+The following analysis focuses on the three most prevalent content categories: Advocacy and recovery, Entertainment, and Educational content - which together comprise 75.2% of coded videos.
+Full cross-tabulations for all categories are provided in Table A.2 in the Appendix.
+Advocacy, Recovery, and Support Content: Videos coded primarily as advocacy and recovery content (33.9%, $n=119$ videos), the largest thematic group in our manual coding, showed strong alignment with two major hashtag communities (see Table A.2 in Appendix):
+  • Awareness & Advocacy (57.2%, $n=99$ videos): Recovery-focused videos frequently used hashtags like #harmreduction and #wedorecover.
+Notably, while recovery-focused hashtags were more commonly used overall, content tagged with harm reduction messaging (particularly #harmreductionsaveslives) generated substantially higher user engagement across all metrics - views, likes, comments, and shares (Figure 3).
+  • Health Conditions (56.8%, $n=100$ videos): Videos discussing recovery often included medical or health-focused tags like #addiction, #chronicpain, and #pain (Figure 4)
+Content in this group was typically centered on active or past efforts to recover from substance use, framed through cautionary narratives or expressions of care regarding the potential harms of substance misuse.
+Figure 4. Sub-networks of Awareness & Advocacy (Pink) and Health Conditions (Blue) to visualize strong overlap between hashtag usage.
+Entertainment and Platform Engagement Entertainment and humor-focused content (28.2%, $n=99$ videos) as well as videos that documented using substances for engagement (4%, $n=14$ videos) also bridged multiple communities.
+These videos predominantly contained hashtags from the following communities:
+  • Platform (40.6%, $n=82$ videos): Entertainment videos heavily utilized platform-specific tags like #fyp and #pingtok
+  • Humor (78.9%, $n=60$ videos): Videos using #addictionhumor and similar tags often presented substance use through comedy or satire
+  • Substance Effects (53.8%, $n=35$ videos): Videos that were satirical or relatable often documented an effect of substance use (e.g., what it’s like being #high or #stoned).
+The subjects in these videos may not necessarily be intoxicated, but could be enacting or portraying the effects.
+While substance use and misuse remained a topic in these videos, humor served as the dominant framing.
+A large portion of videos in this group were meme-based, in which non–drug-related content was reframed through captions, filters, or contextual cues, or drug-related content was presented primarily for comedic effect rather than instructional or cautionary purposes.
+Figure 5. Distribution of video topics across three levels of substance visibility (no substances shown, substances shown but not used, and active substance use) in TikTok videos.
+Educational and Informational Content Videos coded as informational (13.1%, $n=46$ videos) showed connections to multiple professional and health-focused communities:
+  • Health Conditions (11.4%, $n=20$ videos): Medical information and health impacts
+  • Awareness & Advocacy (16.8%, $n=29$ videos): Harm reduction education
+  • Other Substances (24.6%, $n=16$ videos) :
+  • Occupation (19.2%, $n=10$ videos): Content from healthcare workers and other professions
+Content in this group was predominantly clinical or informational in tone.
+Humor was used sparingly, if at all.
+Most videos relied on detailed captions or longer-form explanations, often delivered by health professionals, recovery advocates, or organizations, many of which exhibited higher production value.
+Cross-Cutting Communities.
 While advocacy, entertainment, and educational content showed strong alignment with specific hashtag communities, certain clusters served as contextual bridges across all content types.
-Hashtags from the Commonly Misused Substances (111 videos, 31.6%), Identity & Community (n = 95 videos, 27.1 %), and Occupation (53 videos, 15.1 %) clusters appeared across all seven content categories, suggesting their role in providing context rather than signaling specific content types.
-For instance, videos tagged with “#opiates” spanned multiple content categories: healthcare workers explaining naloxone administration (Recovery/Advocacy), recovery nurses humorously discussing their reliance on caffeine (Satirical), and medical professionals explaining medication effects (Informational).
-Similarly, occupation-related hashtags like “#nurse” or “#doctor” and identity-based tags like “#disabled” or “#veteran” appeared across all content.
-Consumption Method hashtags like “#vape” or “#smoke” (101 videos, 28.8% of videos) served a more specific bridging function, describing substance use behaviors across content types.
-RQ3: Effects of Substance Visibility
-Given these patterns in how content creators use hashtags to frame and contextualize their videos, we next examined a key aspect of substance-related content: the inclusion of substances or their related paraphernalia themselves in the videos with respect to context and engagement.
-This analysis revealed distinct patterns in content distribution across three categories: no substances shown (n=260), substances shown but not used (n=68), and active substance use (n=23).
-Videos showing active substance use were predominantly focused on documentation for social media (approximately 40% of videos, Figure 5), with a stronger tendency towards lifestyle or vlog-style content.
-In contrast, recovery advocacy content was most prevalent in videos without substance visibility, while promotional content appeared most frequently in videos that showed but did not depict active use of substances (e.g., videos promoting nicotine products).
+Hashtags from the Commonly Misused Substances (31.6%, $n=111$ videos), Identity & Community (27.1 %, $n=95$ videos), and Occupation (15.1 %, $n=53$ videos) clusters appeared across all seven content categories, suggesting their role in providing context rather than signaling specific content types.
+For instance, videos tagged with #opiates spanned multiple content categories: healthcare workers explaining naloxone administration (Recovery/Advocacy), recovery nurses humorously discussing their reliance on caffeine (Satirical), and medical professionals explaining medication effects (Informational).
+Similarly, occupation-related hashtags like #nurse or #doctor and identity-based tags like #disabled or #veteran appeared across all content.
+Consumption Method hashtags like #vape or #smoke (28.8%, $n=101$ videos) served a more specific bridging function, describing substance use behaviors across content types.
+4.3. RQ3: Effects of Substance Visibility
+Given the observed patterns in how content creators use hashtags to frame and contextualize their videos, we next examined how the visibility of substances or related paraphernalia within videos relates to content type and engagement patterns.
+Videos were categorized into three groups based on substance presence: no substances shown ($n=260$), substances shown but not actively used ($n=68$), and active substance use ($n=23$).
+Videos showing active substance use were predominantly focused on documentation for social media (39.1% of such videos; Figure 5), with a stronger tendency towards lifestyle or vlog-style content.
+In contrast, recovery advocacy content was most prevalent in videos without substance visibility (38.8%), while promotional content appeared most frequently in videos that showed but did not depict active use of substances (29.4%; e.g., videos promoting nicotine products).
 Within videos showing substances, content predominantly featured alcohol, nicotine, or cannabis products, with only one video collage containing “hard drugs” such as Percocet, Adderall, and Xanax.
-A Kruskal-Wallis H (Kruskal and Wallis 1952) test was performed to compare the means in engagement metrics (views, likes, comments, and shares) between the substance presence categories.
-Notably, there were no statistically significant differences between groups.
-This suggests that user engagement on TikTok is driven more by content type and presentation than by the mere presence or use of substances.
-Discussion
-Social media’s impact on substance use behaviors has been a growing concern, with platforms often criticized for normalizing or promoting dangerous behaviors (Curtis et al.
-2018; Moreno and Whitehill 2014; Lobstein et al.
-2017; Rutherford et al.
-2023; Romer and Moreno 2017).
-Our analysis of TikTok, however, revealed a more nuanced reality where recovery and harm reduction messaging dominate substance-related discourse.
-We found evidence of TikTok emerging as a significant platform for recovery support while maintaining its core identity as an entertainment platform.
-The Power of Digital Recovery Narratives The most striking finding is the prevalence of recovery-focused content, which typically feature first-person narratives celebrating sobriety milestones or sharing recovery journeys.
-As Slade et al.
-(2021) note, such “digital recovery narratives” can increase knowledge of evidence-based treatment options, normalize addiction experiences, and encourage help-seeking behaviors.
-The impact of these narratives appears to be bidirectional - benefiting creators and viewers.
-For creators, sharing recovery stories can reinforce recovery-focused, collective identities and provide accountability through community support (Beck et al.
-2023).
-For viewers, these narratives offer realistic models of recovery success, reduce feelings of isolation, and demonstrate that recovery is possible across different circumstances (Jacobson 2001; Kaiser et al.
-2020; Corvini, Golomski, and Burns 2024).
-Balancing Entertainment and Support in Health Communication The significant presence of humorous content (28.2%) alongside serious recovery narratives reflects TikTok’s unique ability to facilitate health communication through multiple approaches.
-Rather than undermining serious messages, humor appears to serve as an additional entry point for engaging with recovery-related content.
-The use of humor in substance-related content may serve important psychological functions, as research has shown that comedy can help individuals process difficult experiences (Martin 2002, 2003) and reduce stigma around mental health challenges (Corrigan et al.
-2014).
-In the context of substance use and recovery on TikTok, videos that integrate humor appear to make the challenges around substance use more approachable while facilitating community connection through shared experiences.
-This balance between entertainment and support may be particularly effective for reaching younger audiences.
-However, this dual nature also presents challenges.
-While most humorous content avoids glorifying substance use some videos may inadvertently normalize concerning behaviors through comedy.
-This tension between engagement and responsible messaging represents a key challenge for platforms hosting health-related content.
-Network Structure Facilitates Community Building The network analysis revealed a highly interconnected structure with small-world properties, enabling both focused discussions within communities and efficient information flow between them.
-Hashtags serve multiple functions: some act as content summarizers (e.g., “#harmreduction”, “#recovery”), while others serve as contextualizers (e.g., identity-based tags, occupation tags).
-Within this network structure, we observed the presence of algospeak in the form of coded or evasive hashtags (e.g., “#drvg”, “#drvgzx”, “#dr00gtiktok”) that are designed to evade content moderation.
-Our content analysis revealed that these coded hashtags were predominantly associated with satirical and relatable content, with a small subset appearing in recovery advocacy videos.
-This pattern suggests these alternative spellings may serve less to promote substance use and more to facilitate humor-based discussions and shared experiences around substance use, though this practice could complicate platforms’ efforts to effectively moderate substance-related content (Klug, Steen, and Yurechko 2023; Levine 2022).
-This dual functionality helps explain why content tagged with harm reduction and recovery-focused hashtags received substantially higher engagement despite appearing less frequently in our sample.
-Looking Forward: Implications for Social Media Health Communication
-Our findings have important implications for how social media platforms approach substance-related content, not just on TikTok but across all platforms featuring short-form video content like Instagram Reels and YouTube Shorts.
-The success of recovery narratives suggests that brief, personal videos may be particularly effective for health communication.
-Rather than focusing solely on restricting content, platforms might consider ways to actively promote recovery and harm reduction content, given its demonstrated appeal to users.
-The emergence of TikTok as a recovery support space (Russell et al.
-2022) offers valuable lessons for content moderation and platform design.
-While current approaches appear to effectively limit harmful content while allowing space for support and advocacy, refinements could address the challenges of moderating humor-based content.
-As short-form video becomes an increasingly dominant mode of social media communication, particularly among younger users, understanding how to foster supportive recovery communities while mitigating potential risks becomes important.
-Several areas warrant further research.
-While we found evidence of TikTok serving as a positive space for recovery support, there is a pressing need to understand the real-world impacts of this online discourse.
-Future studies should examine how exposure to recovery content affects help-seeking behaviors and treatment outcomes.
-Additionally, researchers should investigate potential risks and benefits for people who use drugs who engage with or create TikTok content.
-This is particularly important for content creators who publicly identify as being in active recovery, as sharing personal narratives may have complex psychological and social implications that are not yet well understood (Rennick-Egglestone et al.
-).
-Additionally, investigating the intersection of online and offline support networks could provide insights into how digital recovery communities complement traditional recovery support services, optimizing online and in-person interventions.
-This work would be particularly valuable for reaching younger populations, who may be more likely to seek initial support through social media platforms.
-Conclusion
-This study explored the use of over 2.3K substance-related hashtags used on the popular short-form social media platform, TikTok.
-Through a semi-automatic community detection process, we identified 16 distinct hashtag communities, including Advocacy and Awareness, Health Conditions, and Humor.
-Our social network analysis revealed a highly interconnected small-world network (clustering coefficient 0.52, average path length 2.20) with influential hashtags like “#addiction”, “#recovery”, and “#sober” serving as bridges between communities.
-Through manual coding of 351 representative videos, we found that Recovery Advocacy content dominated (33.9% of videos), followed by Satirical content (28.2%), suggesting TikTok serves primarily as a platform for support rather than substance promotion.
-This is further supported by our finding that substances were present in only 91 videos (26%), with active use shown in just 23 videos.
-The alignment between hashtag communities and video content, particularly in recovery-focused videos, indicates organic community formation rather than attempts to evade content moderation.
-These findings have important implications for platform policies and interventions - while the prevalence of recovery-focused content suggests TikTok may serve as a valuable space for support and advocacy, the significant presence of humorous content alongside serious advocacy highlights the need for nuanced content moderation approaches.
-Future research should examine how these digital recovery communities complement traditional recovery support services and their real-world impacts on help-seeking behaviors and treatment outcomes, particularly among youth.
-Limitations
-Several limitations of our study warrant discussion.
+Kruskal-Wallis H tests (kruskal_use_1952) compared engagement metrics (views, likes, comments, and shares) across substance visibility categories (Table 4).
+Results indicated a statistically significant difference in likes ($H=6.97$, $p=.031$), with videos depicting active substance use showing higher median likes (Mdn =13,000=13{,}000) compared to videos showing substances without use (Mdn =982=982) or no substances (Mdn =1,028=1{,}028).
+Differences in views ($p=.087$), comments ($p=.171$), and shares ($p=.093$) did not reach statistical significance at $\alpha=.05$.
+However, descriptive patterns consistently showed higher median engagement for videos depicting active substance use across all metrics - notably, median views were over five times higher (79,100 vs.
+14,900) and median comments nearly four times higher (150 vs.
+39) compared to videos without substance visibility.
+Given the small sample size for active use videos ($n=23$), these results should be interpreted cautiously.
+Nevertheless, the findings suggest that the nature of substance depiction may influence audience engagement, with active use content attracting greater interaction despite comprising a small fraction of the overall substance-related content landscape.
+Median (IQR)
+Metric No Showing Using H p
+$(n=260)$ $(n=68)$ $(n=23)$
+Views 14,900 22,150 79,100 4.88 .087
+(1,972–86,775) (968–234,175) (12,450–263,200)
+Likes 1,028 982 13,000 6.97 .031*
+(160–6,146) (43–13,275) (1,222–42,250)
+Comments 39 34 150 3.53 .171
+(7–173) (2–172) (21–253)
+Shares 20 24 112 4.76 .093
+(2–198) (1–458) (28–350)
+Table 4. Engagement metrics by substance visibility category. Kruskal-Wallis H tests compared engagement across visibility levels. IQR = Interquartile Range. *$p<.05$
+5. Discussion
+Our findings present a critical case study of a socio-technical ecosystem where user agency and platform governance collide.
+While platforms like TikTok are often critiqued for amplifying harmful behaviors, including eating disorders (Bickham2024HiddenInPlainSight), self-harm content (lookingbill2024always), and dangerous viral challenges (Curtis et al., 2018; rutherford_turnttrending_2023), our analysis of its algorithmically surfaced content reveals a more complex reality.
+These insights emerge from a mixed-methods analysis that combines hashtag network structure with qualitative coding of algorithmically surfaced video content, allowing us to examine not only what content exists but what is made visible at scale.
+We find that, far from being a simple vector for promotion, the visible discourse on substance use is dominated by digital recovery narratives and harm reduction advocacy.
+This suggests that a combination of community-driven needs, user-led identity work, and active platform moderation has cultivated a powerful, large-scale recovery support space.
+This creates a fundamental tension: while the platform’s recommendation algorithm has documented harms, this same algorithmic infrastructure enables the formation of robust recovery support communities.
+This “algorithmic ambivalence” (phillips2018ambivalence) poses critical challenges for platform governance, as interventions designed to mitigate harm risk dismantling these beneficial support structures.
+5.1. The Politics of Visibility in Digital Recovery Narratives
+The dominance of recovery-focused content (33.9% of videos) suggests that TikTok functions as a venue for what Slade et al.
+(2021) term “digital recovery narratives.” These narratives serve a dual function.
+For creators, the act of public declaration serves as a form of ‘stigma management’ (goffman2009stigma), where the platform’s affordances for visibility are used to reframe a marginalized identity into one of strength and advocacy.
+For viewers, these narratives offer realistic models of recovery success, offer avenues for connection (Phelan2025SocialMediaRecovery), and provide ‘social proof’ (cialdini2004social) of sobriety, countering fatalistic views of addiction (corvini_impact_2024).
+This mirrors findings by li_explaining_2025, who identified that ideological homogeneity and central agendas are key to sustained participation in online movements; here, the ‘movement’ is recovery, and the shared identity is the driver of community.
+However, this visibility raises questions regarding representation and power.
+While our findings demonstrate that recovery advocacy content achieves algorithmic prominence, understanding which recovery narratives are amplified, and which are marginalized, requires further investigation.
+The Phelan2025SocialMediaRecovery finding that individuals in SUD recovery actively use platforms like TikTok for support underscores the stakes of these visibility dynamics (Phelan2025SocialMediaRecovery).
+The absence of active use content in our sample does not necessarily imply its absence from the platform, but rather its algorithmic suppression or relegation to private spaces (gillespie2018custodians).
+This reflects a form of “uneven data shadows” (shelton2014mapping), where certain users and narratives, potentially including those in active addiction, those from marginalized communities, or those whose experiences don’t align with platform-favored content formats, remain algorithmically invisible.
+Consequently, claims about substance use on TikTok must be qualified as claims about what the platform’s governance systems allow to be widely visible.
+5.2. Humor as Boundary Work and Stigma Management
+The significant presence of humorous content (28.2%) alongside serious recovery narratives reflects a distinct form of health communication that we conceptualize as “boundary work” (gieryn1983boundary).
+In this context, humor helps users navigate the boundaries of platform acceptability and social stigma.
+By framing substance use experiences through satire and memes (shifman2014memes), creators may discuss addiction without the immediate vulnerability of serious or sensitive disclosure (andalibi2017sensitive), potentially lowering the barrier to entry for community participation.
+This finding aligns with rifat_combating_2024’s (rifat_combating_2024) observation that community-specific cultural practices—in their case, religious traditions, and in ours, vernacular humor—are essential for mediating online conflict and harm.
+While research has shown that comedy can help individuals process difficult experiences (Martin, 2002), the prevalence of humor in this domain can also introduce risks of minimization.
+The algorithmic favoring of high-engagement humorous content (bucher2017algorithmic) could pressure creators to frame serious health crises comedically to achieve visibility, potentially distorting the perceived severity of substance use disorders for outside observers (chou2018addressing).
+This tension between therapeutic humor and potential minimization represents a key challenge for platforms hosting health-related content, suggesting that engagement metrics alone are insufficient proxies for the quality of health support.
+5.3. Network Structure and the Cooperative Negotiation of Space
+Our network analysis challenges the prevailing narrative of “adversarial algospeak,” where users are assumed to be constantly evading moderation through coded language (chancellor2016thyghgapp; Klug et al., 2023).
+While we observed some coded hashtags (e.g., #dr00gtiktok), the strong alignment between content types and straightforward hashtag communities (e.g., #harmreduction, #wedorecover) suggests that users are largely creating discourse within the bounds of platform affordances (boyd2007social), rather than purely against them.
+This indicates a cooperative, albeit fragile, negotiation of space where the community has successfully signaled the legitimacy of recovery discourse to the platform’s moderation systems, whether through community signaling, policy evolution, or emergent algorithmic behavior.
+Structurally, we found that recovery-focused hashtags function as essential “bridge nodes” (burt1992structural) in the network, connecting otherwise siloed communities.
+This small-world (watts_collective_1998) property enables efficient information flow between disparate groups, such as those focused on specific substances and those focused on sobriety.
+This structural insight is crucial for governance: blunt-force moderation strategies, such as banning specific keywords, could sever these bridges, fragmenting the network and isolating users in riskier sub-communities without pathways to support (chandrasekharan2017you).
+5.4. Implications for Harm-Aware Platform Design
+Our findings suggest that platform governance must move beyond a binary of “remove” or “allow” (gillespie2018custodians) toward a strategy of context-sensitive amplification.
+We propose several concrete design directions:
+1- Leverage network topology for harm reduction pathways.
+Our identification of recovery-focused hashtags as bridge nodes (#addiction, #recovery, #harmreduction) reveals structural leverage points for intervention.
+Rather than simply suppressing substance-related content, recommendation algorithms (resnick1997recommender) could proactively surface content from these bridge communities to users engaging with substance-specific or active-use clusters.
+For example, users repeatedly engaging with #opiates or #pingtok content could receive algorithmically-inserted suggestions for #harmreduction or #wedorecover content, creating pathways from substance-specific communities toward support resources.This approach aligns with public health principles of “meeting people where they are”(marlatt2012harm) while leveraging the small-world network structure (watts_collective_1998) we identified.
+2- Differentiate content types in moderation systems.
+Our codebook (Table 3) demonstrates that recovery advocacy, harm reduction information, active use depiction, and community humor are distinguishable categories.Current moderation systems often treat all substance-related content uniformly (gillespie2018custodians), but context-sensitive classification—potentially incorporating community input (jhaver_does_2019) or subject-matter expertise—could enable more nuanced interventions.
+Content depicting active use without educational framing might be age-gated or de-amplified, while recovery narratives and harm reduction information could be preserved or even amplified.
+However, such systems risk creating new forms of surveillance (eubanks2025automating) and must be designed with strong privacy protections and appeal mechanisms (klonick2018new).
+3- Design for semi-private peer support.
+The current “public-by-default” architecture of TikTok (INPLP_TikTok_2024) creates tensions between visibility (for community-building) and privacy (to avoid stigma (goffman2009stigma)).
+Platforms could offer middle-ground affordances: opt-in recovery communities visible to interested users but not broadcast to followers, pseudonymous accounts specifically for health disclosure, or fine-grained content controls allowing users to share recovery content with support networks while keeping it separate from professional/family audiences (marwick2011tweet).
+Such features would need careful design to avoid inadvertently marking recovery content as shameful.
+4- Involve affected communities in governance.
+Our finding that hashtag communities show organic alignment rather than adversarial evasion suggests potential for collaborative governance (suzor2019we).
+Rather than purely top-down moderation, platforms could establish advisory structures including people with lived experience, harm reduction advocates, and public health professionals.
+Such approaches could help navigate difficult trade-offs between protecting users and preserving support communities, while ensuring governance decisions don’t reproduce existing power inequities in who gets to define “appropriate” recovery discourse (noble2018algorithms).
+Critical to all these recommendations is recognition that the same affordances enabling harm also enable help.
+Well-intentioned changes risk unintended consequences (friedman2013value): special treatment could stigmatize recovery content, surveillance infrastructures built for safety could be weaponized against marginalized users (eubanks2025automating), and clinical norms imposed on peer communities could undermine their organic character (solomon2004peer).
+Effective governance requires explicitly designing for this duality rather than seeking one-sided solutions.
+5.5. The Politics of Algorithmic Invisibility
+Our methodology, analyzing algorithmically-served, publicly visible content, captures what TikTok’s governance systems allow to surface (gillespie2018custodians), not the totality of substance-related discourse.
+This is not merely a limitation but a finding: the dominance of recovery content in our sample reflects active platform curation (gillespie2018custodians) through moderation policies and algorithmic amplification.
+What remains invisible reveals as much about platform governance as what is visible (shelton2014mapping).
+Content in private accounts, shadow-banned videos (gillespie2010politics), age-gated material, and algorithmically suppressed posts likely contain different narratives—potentially including active use content, harm reduction information deemed “controversial,” or experiences from marginalized users whose content may not align with algorithmically-favored formats or whose access to platform visibility is constrained.
+The absence of active use depiction in our sample (only 6.5%) does not prove TikTok is free of such content (bishop2018anxiety); rather, it demonstrates that TikTok’s recommendation system successfully prevents this content from wide circulation.
+This raises critical questions: Should platforms amplify certain recovery modalities over harm reduction approaches?
+Who decides what constitutes “appropriate” substance use discourse?
+Our data suggests TikTok has already made these decisions, whether through explicit policy or emergent algorithmic behavior (gillespie2010politics), in ways that privilege certain recovery modalities and demographic groups.
+The most vulnerable users—those in active addiction, those with intersecting marginalized identities (crenshaw1991mapping), those in acute crisis—may be least likely to appear in algorithmically-surfaced public content (noble2018algorithms), raising concerns about whose needs platform governance serves.
+6. Limitations
+Our findings must be interpreted within several important constraints.
 First, our findings may not fully represent the entire TikTok ecosystem, as they are influenced by limitations in data availability and collection processes.
@@
 Additionally, some videos contained duplicate hashtags, which could affect network metrics, and less popular hashtags may be underrepresented in our sample despite efforts to ensure broad coverage.
-These limitations highlight the challenges of studying substance-related content on social media platforms and suggest the need for complementary research approaches to fully understand these communities.
-Appendix A Appendix
-Appendix B Full Hashtag Network with Communities
-Figure 6 shows a visualization of the complete hashtag co-occurrence network, containing 2,333 nodes (hashtags) and 46,990 edges.
+Lastly, our study analyzed content characteristics rather than user impacts.
+We cannot make claims about whether exposure to recovery content affects help-seeking behaviors, treatment outcomes, or relapse risks.
+The relationship between what people see on TikTok and their subsequent actions remains an open question.
+These limitations underscore a fundamental point: platform research always captures a mediated reality shaped by technical systems, corporate governance, and social dynamics.
+Claims about “what happens on TikTok” must be qualified as “what TikTok’s governance systems allow to be widely visible.”
+7. Conclusion and Future Direction
+This study challenges narratives about social media platforms and substance use.
+Through analysis of 2,333 hashtags and 351 videos, we found that TikTok’s algorithmically surfaced, publicly visible substance use content is dominated by recovery advocacy (33.9%) and humor-based coping (28.2%), not glamorization or promotion.
+Our network analysis revealed a highly interconnected small-world structure (clustering coefficient 0.52, average path length 2.20) where recovery-focused hashtags like #addiction, #recovery, and #sober function as critical bridges between thematic communities.
+Through manual coding of 351 representative videos, we found that Recovery Advocacy content dominated (33.9% of videos), followed by Satirical content (28.2%), while direct substance depiction appeared in only 26% of videos, with active use shown in just 6.5%—suggesting that despite documented harms from algorithmic recommendation systems, these same systems can enable valuable peer support communities around stigmatized health conditions.
+However, our findings also reveal critical tensions inherent in platform governance.
+The dominance of recovery content in our sample reflects TikTok’s active curation through moderation policies and algorithmic amplification—raising questions about whose narratives are privileged and what remains invisible.
+The alignment between hashtag communities and content types indicates organic community formation rather than adversarial evasion, suggesting users have successfully negotiated space within platform affordances.
+Yet this cooperative relationship remains fragile: changes in moderation policy, algorithmic updates, or increased regulatory pressure could easily disrupt these communities, pushing discourse underground or fragmenting support networks.
+Our findings open several critical avenues for future research.
+While we found evidence of TikTok serving as a positive space for recovery support, there is a pressing need to understand the real-world impacts of this online discourse.
+Future studies should examine how exposure to recovery content affects help-seeking behaviors and treatment outcomes.
+Additionally, researchers should investigate potential risks and benefits for people who use drugs who engage with or create TikTok content.
+This is particularly important for content creators who publicly identify as being in active recovery, as sharing personal narratives may have complex psychological and social implications that are not yet well understood (rennick-egglestoneMentalHealthRecovery2019).
+Additionally, investigating the intersection of online and offline support networks could provide insights into how digital recovery communities complement traditional recovery support services, optimizing online and in-person interventions.
+This work would be particularly valuable for reaching younger populations, who may be more likely to seek initial support through social media platforms.
+8. Ethics Statement
+This research was deemed exempt by our institution’s IRB as it analyzes publicly available data with no interaction with human subjects.
+In our analysis of TikTok videos, we followed established practices for ethical social media research (fiesler2020no) and we took several steps to protect user privacy: (1) all analysis was conducted on aggregated data, (2) no usernames or personally identifiable information are included in our results, and (3) our public dataset will only contain video URLs and associated metadata, allowing other researchers to hydrate the data while respecting user privacy.
+All collected and annotated data was stored on secure servers approved by our institution.
+These steps ensure that potential negative outcomes due to the use of these data are minimized while maintaining the reproducibility of our research.
+The dataset and associated code will be made available upon acceptance.
+Appendix A.1 Full Hashtag Network with Communities
+Figure A.1 shows a visualization of the complete hashtag co-occurrence network, containing 2,333 nodes (hashtags) and 46,990 edges.
 The nodes are colored according to their community membership as determined by our semi-automated community detection process.
 The network layout was generated using a Fruchterman-Reingold algorithm, which positions nodes with stronger connections closer together.
-The central position of recovery-focused hashtags like “#addiction”, “#recovery”, and “#harmreduction” is clearly visible, demonstrating their role as bridges between communities.
+The central position of recovery-focused hashtags like #addiction, #recovery, and #harmreduction is clearly visible, demonstrating their role as bridges between communities.
 Node sizes are scaled according to their degree centrality, highlighting the most connected hashtags in the network.
 This visualization helps illustrate both the small-world properties of the network (clustering coefficient 0.52, average path length 2.20) and the organic formation of thematic communities around substance-related discourse on TikTok.
-Figure 6: The overall network of 2,333 hashtags related to substance use with colors representing each community as shown in the legend
-Appendix C Network Centrality Measures per Community
-Table C presents detailed network statistics and centrality measures for both the complete network and each of the 17 identified communities including Misc.
+Figure A.1. The overall network of 2,333 hashtags related to substance use with colors representing each community as shown in the legend
+Appendix A.2 Network Centrality Measures per Community
+Table A.1 presents detailed network statistics and centrality measures for both the complete network and each of the 17 identified communities including Misc.
 For each community, we report basic network metrics (number of nodes, edges, average degree, weighted degree, clustering coefficient, and average shortest path length) as well as the top 5 most central nodes according to four different centrality measures (betweenness, closeness, degree, and eigenvector centrality).
 The whole network shows high connectivity (average degree 40.28) and clustering (0.52), with core hashtags like “addiction”, “recovery”, and “sober” consistently ranking highly across all centrality measures.
+Network #N #E Avg D Avg WD Avg C Avg SP Avg B Betweenness Closeness Degree Eigenvector
+Whole Network 2333 46990 40.28 513.84 0.52 2.20 0.0005 addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction
+Awareness & Advocacy 279 2164 15.51 334.30 0.89 1.98 0.0010 harmreduction, wedorecover, harmreductionsaveslives, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, narcansaveslives
+Health Conditions 277 1522 10.99 224.09 0.85 2.02 0.0012 addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, detox
+Emotions & Feelings 172 305 3.55 20.65 0.31 2.39 0.0002 peace, onedayatatime, friend, dream, serenity peace, onedayatatime, dream, serenity, friend peace, onedayatatime, dream, friend, serenity peace, onedayatatime, dream, serenity, friend
+Platform 183 146 1.60 39.58 0.12 2.44 0.0004 pingtok, dr00gtiktok, pingertok, chat, fyp pingtok, dr00gtiktok, pingertok, fyp, fyp (shi) pingtok, dr00gtiktok, pingertok, chat, stonertokfyp pingtok, dr00gtiktok, pingertok, chat, fyp
+Identity & Community 155 97 1.25 8.22 0.09 3.21 0.0001 cleanandsober, na, pwud, 30s, women pwud, na, cleanandsober, women, lgbt cleanandsober, pwud, na, 30s, mujeres cleanandsober, pwud, na, lgbtqia, calisober
+Commonly Misused Substances 152 553 7.28 60.47 0.38 2.47 0.0007 opioids, opiates, coke, molly, fent opioids, coke, opiates, molly, uppers opioids, opiates, coke, molly, fent opioids, coke, opiates, fent, molly
+Location 116 19 0.33 1.26 0.00 1.80 ¡0.0001 chinatown, paradise, yucatan, arnolds, filmteyvatislands chinatown, paradise, india, china, texas chinatown, paradise, yucatan, arnolds, india chinatown, india, china, texas, egypt
+Cannabis 51 229 8.98 39.25 0.55 2.04 0.0006 maryjane, weed, cannabis, skunk, herb maryjane, weed, cannabis, herb, grass maryjane, weed, cannabis, herb, grass maryjane, cannabis, weed, herb, grass
+Alcohol 43 92 4.28 56.19 0.41 1.79 0.0006 drunk, tipsy, hammered, shitfaced, bars drunk, hammered, tipsy, shitfaced, bars drunk, hammered, tipsy, shitfaced, bars drunk, hammered, tipsy, bars, shitfaced
+Cognitive Enhancement 24 38 3.17 45.00 0.46 1.64 0.0001 nootropics, modafinil, smartdrugs, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus
+Substance Effects 41 33 1.61 23.76 0.16 2.42 0.0005 high, wasted, energy, blackout, clarity high, wasted, lit, crash, shadowpeople high, wasted, blackout, lit, downers high, wasted, lit, downers, crash
+Consumption Method 41 62 3.02 89.80 0.30 1.86 0.0017 smoke, vape, smoking, vaper, vaping smoke, vaper, vape, vaping, vaped smoke, vaper, vape, vaping, vaped smoke, vaper, vaping, vaped, vape
+Tobacco Nicotine 31 115 7.42 68.90 0.70 1.70 0.0004 ecigs, ecig, cigarette, nicotine, juul ecigs, cigarette, ecig, nicotine, juul ecigs, cigarette, ecig, nicotine, juul cigarette, ecig, ecigs, nicotine, juul
+Other Substances 70 130 3.71 50.11 0.27 2.69 0.0012 narcan, snow, psychedelictok, naloxone, naltrexone narcan, psychedelictok, snow, , shroooms psychedelictok, snow, narcan, mushrooms, naloxone psychedelictok, mushrooms, salvia, mescaline, magicmushroomsadventures
+Humor 56 19 0.68 28.64 0.00 1.90 0.0002 addictionhumor, prank, pingtok⚫️ ⚫️, standupcomedy, satire addictionhumor, jokes, comedia, soberhumor, comediahumor addictionhumor, jokes, comedia, soberhumor, comediahumor addictionhumor, jokes, comedia, soberhumor, comediahumor
+Occupation 73 5 0.14 0.93 0.00 1.50 ¡0.0001 copilots, truckdrivers, actress, nursing, outreachworker copilots, truckdrivers, aviation, construction, creator copilots, truckdrivers, aviation, work, construction copilots, aviation, construction, creator, truckdrivers
+Misc 565 2367 8.38 139.01 0.29 2.46 0.0002 cat, fire, rock, g, pink cat, fire, g, rock, pink cat, fire, g, rock, x cat, fire, g, x, rock
+Table A.1.
+Centrality measures for the whole network and each community.
+#N=Number of Nodes, #E=Number of Edges, Avg D=Average Degree, Avg WD=Average Weighted Degree, Avg C=Average Clustering, Avg SP=Average Shortest Path, Avg B=Average Betweenness Centrality.
+The top five nodes are presented per centrality measure.
+Consumption Method exhibits the highest average betweenness (bolded).
+Appendix A.3 Hashtag Community Distribution per Content Topic
+Table A.2 presents a cross-tabulation of hashtag communities and video content topics, allowing us to examine how different communities align with various types of content.
+This analysis reveals notable patterns in how hashtags from different communities are employed across content types.
+Recovery Advocacy content, the most prevalent in our dataset, shows strong representation across multiple hashtag communities, particularly drawing from the Awareness & Advocacy and Health Conditions communities.
+Satirical content, our second most common category, shows interesting patterns of hashtag usage, frequently incorporating hashtags from Platform and Humor communities while also drawing from more serious communities - suggesting complex layering of messaging in humorous content.
+Informational content shows strong alignment with Professional and Health Condition hashtags, while Promotional content primarily uses hashtags from Consumption Method and Tobacco/Nicotine communities.
+This distribution helps validate our community detection approach by demonstrating meaningful alignment between hashtag communities and content types, while also revealing how creators strategically combine hashtags from multiple communities to reach diverse audiences.
 Communities
@@
 Cognitive Enhancement 1 (4.0%) 7 (28.0%) 3 (12.0%) 7 (28.0%) 4 (16.0%) 3 (12.0%) 0 (0.0%) 25
-Table 4: Distribution of communities corresponding with each hashtag used in the 351 analyzed videos per content topic.
-Network #N #E Avg D Avg WD Avg C Avg SP Betweenness Closeness Degree Eigenvector
-Whole Network 2333 46990 40.28 513.84 0.52 2.20 addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction
-Awareness & Advocacy 279 2164 15.51 334.30 0.89 1.98 harmreduction, wedorecover, harmreductionsaveslives, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, narcansaveslives
-Health Conditions 277 1522 10.99 224.09 0.85 2.02 addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, detox
-Emotions & Feelings 172 305 3.55 20.65 0.31 2.39 peace, onedayatatime, friend, dream, serenity peace, onedayatatime, dream, serenity, friend peace, onedayatatime, dream, friend, serenity peace, onedayatatime, dream, serenity, friend
-Platform 183 146 1.60 39.58 0.12 2.44 pingtok, dr00gtiktok, pingertok, chat, fyp pingtok, dr00gtiktok, pingertok, fyp, fyp:) pingtok, dr00gtiktok, pingertok, chat, stonertokfyp pingtok, dr00gtiktok, pingertok, chat, fyp
-Identity & Community 155 97 1.25 8.22 0.09 3.21 cleanandsober, na, pwud, 30s, women pwud, na, cleanandsober, women, lgbt cleanandsober, pwud, na, 30s, mujeres cleanandsober, pwud, na, lgbtqia, calisober
-Commonly Misused Substances 152 553 7.28 60.47 0.38 2.47 opioids, opiates, coke, molly, fent opioids, coke, opiates, molly, uppers opioids, opiates, coke, molly, fent opioids, coke, opiates, fent, molly
-Location 116 19 0.33 1.26 0.00 1.80 chinatown, paradise, yucatan, arnolds, filmteyvatislands chinatown, paradise, india, china, texas chinatown, paradise, yucatan, arnolds, india chinatown, india, china, texas, egypt
-Cannabis 51 229 8.98 39.25 0.55 2.04 maryjane, weed, cannabis, skunk, herb maryjane, weed, cannabis, herb, grass maryjane, weed, cannabis, herb, grass maryjane, cannabis, weed, herb, grass
-Alcohol 43 92 4.28 56.19 0.41 1.79 drunk, tipsy, hammered, shitfaced, bars drunk, hammered, tipsy, shitfaced, bars drunk, hammered, tipsy, shitfaced, bars drunk, hammered, tipsy, bars, shitfaced
-Cognitive Enhancement 24 38 3.17 45.00 0.46 1.64 nootropics, modafinil, smartdrugs, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus
-Substance Effects 41 33 1.61 23.76 0.16 2.42 high, wasted, energy, blackout, clarity high, wasted, lit, crash, shadowpeople high, wasted, blackout, lit, downers high, wasted, lit, downers, crash
-Consumption Method 41 62 3.02 89.80 0.30 1.86 smoke, vape, smoking, vaper, vaping smoke, vaper, vape, vaping, vaped smoke, vaper, vape, vaping, vaped smoke, vaper, vaping, vaped, vape
-Tobacco Nicotine 31 115 7.42 68.90 0.70 1.70 ecigs, ecig, cigarette, nicotine, juul ecigs, cigarette, ecig, nicotine, juul ecigs, cigarette, ecig, nicotine, juul cigarette, ecig, ecigs, nicotine, juul
-[…]
-[…]
-Occupation 73 5 0.14 0.93 0.00 1.50 copilots, truckdrivers, actress, nursing, outreachworker copilots, truckdrivers, aviation, construction, creator copilots, truckdrivers, aviation, work, construction copilots, aviation, construction, creator, truckdrivers
-Misc 565 2367 8.38 139.01 0.29 2.46 cat, fire, rock, g, pink cat, fire, g, rock, pink cat, fire, g, rock, x cat, fire, g, x, rock
-Table 5: Centrality measures for the whole network and each community.
-#N=Number of Nodes, #E=Number of Edges, Avg D=Average Degree, Avg WD=Average Weighted Degree, Avg C=Average Clustering, Avg SP=Average Shortest Path.
-The top five nodes are presented per centrality measure.
-Appendix D Hashtag Community Distribution per Content Topic
-Table 4 presents a cross-tabulation of hashtag communities and video content topics, allowing us to examine how different communities align with various types of content.
-This analysis reveals notable patterns in how hashtags from different communities are employed across content types.
-Recovery Advocacy content, the most prevalent in our dataset, shows strong representation across multiple hashtag communities, particularly drawing from the Awareness & Advocacy and Health Conditions communities.
-Satirical content, our second most common category, shows interesting patterns of hashtag usage, frequently incorporating hashtags from Platform and Humor communities while also drawing from more serious communities - suggesting complex layering of messaging in humorous content.
-Informational content shows strong alignment with Professional and Health Condition hashtags, while Promotional content primarily uses hashtags from Consumption Method and Tobacco/Nicotine communities.
-This distribution helps validate our community detection approach by demonstrating meaningful alignment between hashtag communities and content types, while also revealing how creators strategically combine hashtags from multiple communities to reach diverse audiences.
+Table A.2. Distribution of communities corresponding with each hashtag used in the 351 analyzed videos per content topic.